### v1

# MS-Mix: Unveiling the Power of Mixup for Multimodal Sentiment Analysis

Abstract.
Multimodal Sentiment Analysis (MSA) aims to identify and interpret human emotions by integrating information from heterogeneous data sources such as text, video, and audio. While deep learning models have advanced in network architecture design, they remain heavily limited by scarce multimodal annotated data. Although Mixup-based augmentation improves generalization in unimodal tasks, its direct application to MSA introduces critical challenges: random mixing often amplifies label ambiguity and semantic inconsistency due to the lack of emotion-aware mixing mechanisms. To overcome these issues, we propose MS-Mix, an adaptive, emotion-sensitive augmentation framework that automatically optimizes sample mixing in multimodal settings. The key components of MS-Mix include: (1) a Sentiment-Aware Sample Selection (SASS) strategy that effectively prevents semantic confusion caused by mixing samples with contradictory emotions. (2) a Sentiment Intensity Guided (SIG) module using multi-head self-attention to compute modality-specific mixing ratios dynamically based on their respective emotional intensities. (3) a Sentiment Alignment Loss (SAL) that aligns the prediction distributions across modalities, and incorporates the Kullback-Leibler-based loss as an additional regularization term to train the emotion intensity predictor and the backbone network jointly. Extensive experiments on three benchmark datasets with six state-of-the-art backbones confirm that MS-Mix consistently outperforms existing methods, establishing a new standard for robust multimodal sentiment augmentation. The source code is available at: https://github.com/HongyuZhu-s/MS-Mix.
Multimodal sentiment analysis, Data augmentation, Regularization
copyright: acmlicensedcopyright: acmlicensedjournal: TOMMjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: XXXXXXX.XXXXXXXccs: Computing methodologies Neural networksccs: Computing methodologies Regularization
1. Introduction
The perception and understanding of human emotions by Artificial Intelligence (AI) is of great significance for technologies such as human-computer interaction (Hibbeln et al., 2017) and multimedia computing(Li et al., 2019). Multimodal Sentiment Analysis (MSA) has emerged as a critical research frontier in this endeavor (Dumpala et al., 2019; Liu et al., 2025; Mao et al., 2022; Zhang et al., 2023; Zhuang et al., 2024; Xiao et al., 2025). MSA aims to integrate and interpret complementary emotional cues from textual, acoustic, and visual modalities to achieve a more robust and accurate understanding of sentiment (Manning et al., 2014). The significance of MSA is underscored by its wide-ranging applications, including but not limited to opinion mining on social media (Chandrasekaran et al., 2021), empathetic chatbot design (Yang et al., 2019), and automated mental health assessment (Singh et al., 2024). By moving beyond unimodal analysis, MSA offers the potential to capture the complex and often incongruent nature of human affective expression, thereby providing a more holistic view of sentiment.
Benefiting from the rapid advances in Deep Learning (DL), the field of MSA has witnessed substantial progress in the development of neural architectures. Techniques such as cross-modal transformers (Tsai et al., 2019) and advanced fusion mechanisms (Zhuang et al., 2024) have demonstrated remarkable capability in modeling inter-modal interactions and extracting discriminative features. Nevertheless, despite these architectural advancements, the performance of data-driven DL models remains fundamentally constrained by the scale and quality of annotated data (Zhang et al., 2018). This data scarcity often results in overfitting and limited generalization ability (Srivastava et al., 2014). Furthermore, the construction of MSA datasets is costly, as it necessitates the collection of reliable human-annotated multimodal sentiment dataset.
Fig. 1. The differences between MS-Mix and traditional mixup methods represented by Manifold Mixup (Zhang et al., 2018). Left: a. The traditional mixup method employs random selection of samples and an offline mixing ratios optimization strategy. b. Backbone. c. MS-Mix. Right: MS-Mix can generate augmented samples that better align with the original distribution on the MOSEI dataset. Where the KL represents the Kullback-Leibler divergence. :
To mitigate the data scarcity issue, the mixup strategy, a data augmentation technique, was introduced to generate virtual training samples via convex interpolations between original data points and their corresponding labels (Zhang et al., 2018). However, existing offline methods, including general domain techniques such as CutMix (Yun et al., 2019) and SaliencyMix (Uddin et al., 2021), as well as multimodal specific approaches like $\mathcal{P}ow$Mix (Georgiou et al., 2024) and MultiMix (Venkataramanan et al., 2023), often depend on randomly paired samples and offline optimization strategies, thereby neglecting the underlying emotional semantics. This may result in the blending of semantically inconsistent samples, e.g., combining a positive and a negative sample can introduce semantic confusion and label mismatch. Moreover, these methods typically employ fixed or uniformly distributed mixing ratios, which fail to adapt to the varied emotional intensities across different modalities. As a consequence, the quality and efficacy of the augmented data are substantially compromised.
To address these limitations, we propose MS-Mix, a novel and adaptive data augmentation framework specifically designed for multimodal sentiment analysis. As shown in Fig. 1, we have visualized the distribution of the generated labels. It can be observed that the label distribution generated by MS-Mix is closer to the real situation. Our method introduces three key innovations to ensure semantically consistent and high-quality sample generation: (1) We design a Sentiment-Aware Sample Selection (SASS) strategy that leverages semantic similarity in the latent space to filter out incompatible pairs, thereby preventing the mixture of samples with contradictory emotions. (2) We introduce a Sentiment Intensity Guided (SIG) mixing module to dynamically determine the mixing ratio across modalities, which leverages a multi-head self-attention mechanism to predict modality-specific mixing ratios conditioned on sentiment salience, thereby enabling more fine-grained and context-aware fusion. (3) We introduce a Sentiment Alignment Loss (SAL) to align the predicted sentiment distribution with the ground-truth labels. The SAL acts as a regularization term, enhancing the backbone network’s discriminative power and improving the consistency of representations from augmented samples. Experimental results conducted on the real-world MSA datasets demonstrate that MS-Mix significantly outperforms existing methods across multiple baselines and backbone architectures.
The main contributions of this paper are summarized as follows:
  • We design an SASS strategy that leverages semantic similarity in the latent space to filter out incompatible sample pairs, effectively preventing mixtures of samples with contradictory emotions.
  • We introduce a SIG mixing module, implemented via a multi-head self-attention mechanism, to dynamically determine modality-specific mixing ratios based on emotional salience.
  • We propose SAL, a Kullback-Leibler (KL) divergence-based regularization that aligns predicted sentiment distributions with ground-truth labels, thereby enhancing both model discrimination and consistency.
  • We conduct extensive experiments on three popular MSA datasets and six backbone architectures, demonstrating that MS-Mix significantly outperforms existing methods.
The remainder of this paper is organized as follows. Section 2 reviews related methods. The proposed MS-Mix method is detailed in Section 3. The experimental results are presented in Section 4, and the concluding remarks are provided in Section 5.
2. Related Work
2.1. Multimodal Sentiment Analysis
Sentiment analysis systems are divided into single-modal systems and multi-modal systems. Compared with single-modal systems, which analyze the sentiment state through a single data source, MSA systems can utilize complementary information from various modalities such as text, video, and audio, thereby providing an effective method for comprehensive sentiment analysis (Geetha et al., 2024). For instance, Cimtay et al. (Cimtay et al., 2020) achieved emotion recognition by extracting data features from Galvanic Skin Response (GSR), Electroencephalogram (EEG), and facial expressions. Prior research in MSA has predominantly advanced through improving two main paradigms: feature fusion strategy-centric (Zadeh et al., 2017; Liu and Shen, 2018; Zadeh et al., 2018a; Tsai et al., 2019; Lv et al., 2021; Zhuang et al., 2024) and feature encoder method-centric (Liang et al., 2021; Han et al., 2021; Yang et al., 2022; Zhang et al., 2023) approaches.
Fusion strategy-centric approaches in MSA primarily focused on designing effective fusion strategies to integrate features from different modalities. Zadeh et al. (Zadeh et al., 2017) introduced the Tensor Fusion Network (TFN), which explicitly models inter-modal interactions through the 3-fold cartesian product. Building on this work, Z Liu et al. (Liu and Shen, 2018) proposed Low-rank Multimodal Fusion (LMF), an efficient approach that enhances computational performance via parallel tensor and weight decomposition. By leveraging modality-specific low-rank factors, LMF achieves effective multimodal fusion with significantly reduced computational cost. MuIT (Tsai et al., 2019) incorporates cross-modal attention, exploring the long-range dependency relationships among cross-modal elements in multimodal datasets, and achieving multimodal fusion on unaligned multimodal datasets. Based on the same motivation and algorithm, Huang et al. (Huang et al., 2020) employed cross-modal attention to learn the long-range dependencies between the visual and audio modalities, achieving better performance than decision-level and feature-level fusion. A recent advancement in MSA is the attempt to decouple the features into shared and unique information. For instance, Li et al. (Li et al., 2023) proposed a Decoupled Multimodal Distillation (DMD) method, which is capable of distilling cross-modal knowledge in the decoupled feature spaces and alleviating the problem of inherent multimodal heterogeneity. The Global Local Modal (GLoMo) (Zhuang et al., 2024) Fusion framework integrates multiple local representations within each modality and effectively combines local and global information.
For feature encoder-centric approaches, the primary objective is to enhance the feature representation of each modality. Yang et al. (Yang et al., 2022) conceptualize multimodal representation learning as a form of domain adaptation, employing adversarial learning to model both modality-invariant and modality-specific subspaces within multimodal fusion. Han et al. (Han et al., 2021) introduced the MMIM framework, which enhances multimodal fusion performance through hierarchical mutual information maximization. To address the challenge of multimodal heterogeneity in MSA, Li et al. (Lin and Hu, 2023) conceptualized the learning process of each modality as a set of sub-tasks and introduced a novel Multi-Task Momentum Distillation (MTMD) method. This approach effectively reduces the discrepancies between modalities and enhances representation consistency. Furthermore, to suppress sentiment-irrelevant and conflicting information across modalities, Zhang et al. (Zhang et al., 2023) proposed the Adaptive Language-guided Multimodal Transformer (ALMT), which learns a unified hypermodality representation guided by language at multiple scales. In recent work, Li et al. (Li et al., 2025) proposed the MM-PEAR-CoT framework, which enhances the reliability of multimodal sentiment analysis by using cross-modal filtering and fusion to suppress irrational reasoning steps generated by large language models.
Despite these advancements, MSA systems remain constrained by the scarcity and high annotation cost of multimodal data, often leading to overfitting and limited generalization (Venkataramanan et al., 2023; Georgiou et al., 2024). This highlights the need for effective data augmentation techniques specifically designed for multimodal emotional data.
2.2. Mixup-based Augmentation
To mitigate data scarcity and enhance generalization, data augmentation techniques—particularly Mixup and its variants—have been extensively adopted in DL (Kim et al., 2020a; Liu et al., 2022; Qin et al., 2024a, b; Jin et al., 2024). The original Mixup method (Zhang et al., 2018) operates in the input space by generating virtual samples through linear interpolation between two randomly selected data points and their corresponding labels. This simple yet effective strategy encourages models to behave linearly across classes and improves robustness. Subsequent extensions such as Manifold Mixup (Verma et al., 2019) generalize the interpolation operation to hidden representations, further enhancing the smoothness of decision boundaries. Similarly, in the latent space, the Mixup-Transformer (Sun et al., 2020) encodes a sentence using a Transformer and then linearly interpolates the representations to generate mixed samples for classification. In recent state-of-the-art approaches (Liu et al., 2022; Qin et al., 2024a, b), mixing strategies are no longer manually designed based on prior knowledge or saliency information but instead adaptively generate mixed samples using learnable mixing ratios and feature representations in an end-to-end manner.
In the multimodal domain, several works have extended Mixup to leverage cross-modal interactions. For instance, MultiMix (Venkataramanan et al., 2023) performs independent Mixup operations within each modality and integrates the augmented representations through late fusion. VLMixer (Wang et al., 2022) integrates cross-modal CutMix (Yun et al., 2019) with contrastive learning to convert uni-modal text inputs into multi-modal representations of text and image, thereby improving instance-level alignment across modalities. To enhance robustness against missing modalities, Lin et al. (Lin and Hu, 2024) proposed Multi-Modal Mixup ($M^{3}$ixup), which extends mixup to both representation and contrastive loss levels for improved cross-modal alignment and dynamics capture. More recently, as an improvement of MultiMix, $\mathcal{P}$owMix (Georgiou et al., 2024) introduced a weight-aware mixing strategy that dynamically modulates mixing coefficients based on the estimated importance of each modality, enabling more flexible and effective multimodal augmentation.
However, most existing methods rely on random sample pairing and offline mixing schemes, which overlook the semantic structure and emotional coherence of multimodal data. This often results in semantically inconsistent mixtures, such as mixing samples with opposing emotions, which introduces label noise and hinders model performance. These limitations underscore the need for more semantically aware and adaptively controlled mixing strategies tailored to multimodal sentiment analysis. Therefore, we propose MS-Mix, a novel mixup method that effectively avoids the mixing of contradictory emotions and adaptively generates discriminative multimodal emotional features.
Fig. 2. An overview of the overall structure of the proposed MS-Mix framework. (a) The SASS strategy computes the emotional semantic distance between samples. (b) The SIG module performs adaptive mixing of sample pairs with similar emotional semantics. (c) The SAL ($\mathcal{L}_{SAL}$) serves as an auxiliary regularization term that promotes alignment between predicted emotional intensity values and ground-truth labels via a KL-based loss. These components work collaboratively to enhance multimodal representation learning. :
3. Methodology
In this section, we first introduce the task definition and the overview of our MS-Mix framework. We then describe the details of each module in MS-Mix.
3.1. Task Definition and Framework Overview
Given a MSA dataset with $n$ samples, $\textbf{X}=(\textbf{x}_{1},\textbf{x}_{2},...,\textbf{x}_{n})$} and $\textbf{Y}=(y_{1},y_{2},...,y_{n})$} are the data and labels. For the $i_{th}$ sample $\textbf{x}_{i}\in\textbf{X}$, where $\textbf{x}_{i}$ comprises data from the text (t), video (v), and audio (a) modalities, denoted as $\{\textbf{x}^{t}_{i},\textbf{x}^{v}_{i},\textbf{x}^{a}_{i}\}$. The goal of the MSA task is to learn a mapping function $\mathcal{F}:\textbf{X}\mapsto\textbf{Y}$ to predict the occurrence of each emotion category.
MS-Mix is a novel data augmentation method in the latent space, acting on the multimodal features output by the encoder, and is applied before feature fusion. As illustrated in Fig.2, the input to MS-Mix consists of multimodal features derived from the model’s encoder outputs. The SASS strategy first identifies feature pairs suitable for mixing, which are then adaptively blended through the SIG module to generate augmented features. Both the original and synthesized features are subsequently fed into the fusion module. Additionally, the SAL serves as an auxiliary regularization term to jointly optimize the entire network and the SIG module.
As a common starting point for many mixup-based augmentation methods, vanilla mixup (Zhang et al., 2018) employs a ratio $\lambda$ to construct a mixed sample ($\hat{\textbf{x}},\hat{y}$) by performing global linear interpolation directly on the sample pair $\{(\textbf{x}_{i},y_{i}),(\textbf{x}_{j},y_{j})\}$:
(1) $\hat{\textbf{x}}=\lambda\cdot\textbf{x}_{i}+(1-\lambda)\cdot\textbf{x}_{j},$
$\hat{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j},$
where the $\lambda$ is sampled from $Beta(\alpha,\alpha)$ distribution. We define the process of Eq.1 as follows: Given the sample mixup function $\mathrm{H}(\cdot)$, the label mixup function $\mathrm{G}(\cdot)$, and a mixing ratio $\lambda$, we can generate the mixed sample ($\hat{\textbf{x}},\hat{y}$) with $\hat{\textbf{x}}=\mathrm{H}(\textbf{x}_{i},\textbf{x}_{j},\lambda)$, $\hat{y}=\mathrm{G}(y_{i},y_{j},\lambda)$.
Mixing of original data can easily lead to sentiment semantic confusion in generated data (Verma et al., 2019). Therefore, for each batch of data, we choose to learn different mixing ratios $\lambda^{m}$ to mix the features $\textbf{Z}^{m}=f_{\theta}(\textbf{X})\in\mathbb{R}^{B\times d^{m}}$ of each modal in the latent space separately (Eq. 2), and calculate the label ratio $\lambda^{L}$ to mix the labels as follows:
(2) $\hat{\textbf{z}}^{m}=\mathrm{H}(f_{\theta}(\textbf{x}_{i}^{m}),f_{\theta}(\textbf{x}_{j}^{m}),\lambda^{m}),\textbf{x}^{m}\in\textbf{X},$
(3) $\hat{y}=\mathrm{G}((y_{i},y_{j}),\bm{\lambda}^{L}),y\in\textbf{Y}.$
where $\hat{\textbf{z}}^{m}$ represents the the mixed features, $m\in\{t,v,a\}$ represents a modality belonging to the set of the three modalities in the dataset, $B$ is the mini-batch size, $d^{m}$ is the hidden space dimension of modality $m$, and $f_{\theta}$ is the backbone.
Finally, the mixed features and the original features will be fused into a whole set for the subsequent feature fusion process.
3.2. Sentiment-Aware Sample Selection
Emotional expressions exhibit significant variation across samples. While random in-batch mixing enhances the feature smoothness, it often ignores semantic consistency in the emotional space. This may lead to the mixing of two semantically opposite samples with strong emotion into a mixed sample with a neutral label and semantic confusion features. Therefore, we proposed the SASS strategy based on the emotional semantic distance, and screened the feature samples within a batch before mixing. Specifically, we use the feature cosine distance between each sample to represent the semantic information similarity (Pennington et al., 2014), and mix the sample pairs with a similarity greater than the threshold $\delta$ within each batch. This ensures that only samples with analogous emotions are combined, a critical factor overlooked by previous methods (Georgiou et al., 2024).
We first performed $L_{2}$ normalization on the features of each modality $\textbf{Z}^{m}$ within the batch to ensure consistent feature scaling across modalities:
(4) $\mathbf{Z}^{m}_{\text{norm}}=\frac{\mathbf{Z}^{m}}{\sqrt{\sum_{i=1}^{d^{m}}(Z_{i}^{m})^{2}}}.$
Then, we obtain the similarity matrix S as:
(5) $\textbf{S}=\sum_{m\in t,v,a}{\mathbf{Z}^{m}_{\text{norm}}}\cdot(\mathbf{Z}^{m}_{\text{norm}})^{\top}/3.$
Since the matrix S is symmetric ($s_{ij}=s_{ji},s\in\textbf{S}$) and the diagonal elements represent the self-similarity, we define the normalized upper triangular matrix $\textbf{S}^{\prime}$ as:
(6) $\textbf{S}^{\prime}=\frac{1}{n(n-1)}\sum_{i\neq j}s_{ij}.$
Then we randomly selected $B$ pairs to mix from the feature pairs $\{(\textbf{z}_{i},y_{i}),(\textbf{z}_{j},y_{j})\},\textbf{z}\in\textbf{Z}$, with a similarity greater than $\delta$ = 0.2. Our SASS strategy employs the similarity threshold to exclude sample pairs with opposite emotions before mixing. A lower threshold effectively prevents semantic confusion in the mixed samples while maximizing data diversity.
3.3. Sentiment Intensity Guided Mixing Module
The mix ratio $\lambda$ is one of the most important hyperparameters in mixup-based methods, used to balance the degree of mixing between two or more samples. $\lambda$ is usually sampled from the $Beta(\alpha,\alpha)$ distribution. However, in recent years, some methods (Kim et al., 2020b; Liu et al., 2022; Qin et al., 2024a) have been able to automatically optimize the mixture proportions based on the sample states, thereby achieving better matching between the data and the label.
To ensure that samples with richer emotional semantics contribute more substantially to the mixing process, we propose the SIG mixing module. Specifically, for each modality, a Multi-Head self-Attention (MHA) (Vaswani et al., 2017) encoder is trained to predict emotional intensity values $\textbf{I}^{m}$. This encoder comprises an MHA layer, a residual connection, layer normalization, and a tanh activation function. The modality-specific emotional intensity predictor is denoted as $f_{\phi}^{m}$ (Eq. 7-9). These predictions are then utilized as weighting factors to adjust the mixing ratios $\lambda^{m}$ during augmentation dynamically.
(7) $head_{i}=\mathrm{Softmax}\left(\frac{\textbf{Z}^{m}\textbf{W}_{i}^{Q}\cdot\textbf{Z}^{m}({\textbf{W}_{i}^{K}})^{\top}}{\sqrt{d_{k}}}\right)\textbf{Z}^{m}\textbf{W}_{i}^{V},$
(8) $\mathrm{MHA}(\textbf{Z}^{m})=\mathrm{LN}(\mathrm{Concat}(head_{1},head_{2},.,head_{h})\textbf{W}^{O}+\textbf{Z}^{m}),$
(9) $\textbf{I}^{m}=f_{\phi}^{m}(\textbf{Z}^{m})=\mathrm{tanh}(\mathrm{GobalPool}(\mathrm{MHA}(\textbf{Z}^{m}))).$
where $\textbf{W}^{\{Q,K,V\}}={\{\textbf{W}^{Q},\textbf{W}^{K},\textbf{W}^{V}\}}$ are learnable weight matrices that map the input features to queries $Q$, keys $K$, and values $V$, $d_{k}$ is the dimension of $K$, LNLN denotes Layer Normalization (Ba et al., 2016), the number of attention heads $h$ is set to 4 or 6, and $\textbf{W}^{O}$ is the output linear transformation matrix.
Based on the intensity of emotions $\textbf{I}^{m}$ in Eq. 9, we perform Min-Max normalization on it, which can further calculate the intermediate mixing weights ${\bm{\omega}}^{m}_{i}$ for each feature pair $\{(\textbf{z}_{i},y_{i}),(\textbf{z}_{j},y_{j})\}$:
(10) $\bm{\omega}_{i}^{m}=\frac{|\textbf{I}_{i}^{m}|-min(\textbf{I}^{m})}{max(\textbf{I}^{m})-min(\textbf{I}^{m})+\epsilon}.$
where $\epsilon$ represents the minimum value that prevents division by zero and is set to $10^{-8}$. Due to the mixed strategy of convex combination, the adaptive mixing ratio $\lambda^{m}_{i,j}$ between the $\textbf{z}_{i}$ and $\textbf{z}_{j}$ is:
(11) $\lambda^{m}_{i,j}=(\frac{\bm{\omega}_{i}^{m}}{\bm{\omega}_{i}^{m}+\bm{\omega}_{j}^{m}}+\lambda_{base})/2,$
where the $\lambda_{base}$ sampled from the $\text{Beta}(\alpha,\alpha)$. Then, we calculate the average of the mixing ratios $\lambda^{m}$ for each modality to obtain the mixing ratio $\lambda^{L}_{i,j}$ of the labels:
(12) $\lambda^{L}_{i,j}=\sum_{m\in t,v,a}\lambda_{i,j}^{m}/3.$
Finally, we can get the mixed feature $\hat{\textbf{z}}^{m}$ and mixed label $\hat{y}$ using Eq. 2-3. Unlike the previous Mixup-based studies (Zhang et al., 2018; Verma et al., 2019; Yun et al., 2019; Uddin et al., 2021), our proposed SIG module adaptively determines the mixing ratios between samples based on sentiment intensity, maintaining end-to-end trainability and allowing for continuous optimization throughout the training process.
3.4. Sentiment Alignment Loss Function
To enhance the accuracy of the emotion intensity predictor, we introduce the SAL as an additional regularization term to align the predicted sentiment distribution with the ground-truth labels. In our task, we seek the mapping from the data $x$ to its label $y$ modeled by network $f_{\theta}$, where the network parameters $\theta$ are optimized by minimizing a loss function $\mathcal{L}_{task}$ using Mean Squared Error (MSE):
(13) $\min_{\theta}\mathcal{L}_{task}(f_{\theta}(\textbf{x}),y).$
Since the emotional labels are continuous values (Zadeh et al., 2016, 2018b; Yu et al., 2020), we consider the mixup regression task. Therefore, we get the mapping between the mixed $\hat{\textbf{x}}$ and $\hat{y}$ by optimizing mixed MSE loss $\mathcal{L}_{mixMSE}$ (Eq.14). By minimizing $\mathcal{L}_{mixMSE}$ (Eq.15), the consistency between the mixed samples and labels is constrained.
(14) $\mathcal{L}_{mixMSE}=\lambda^{L}_{i,j}\cdot MSE(f_{\theta}(\hat{\textbf{x}_{i}}),y_{i})+(1-\lambda^{L}_{i,j})\cdot MSE(f_{\theta}(\hat{\textbf{x}_{j}}),y_{j}).$
[…]
The KL divergence (Van Erven and Harremos, 2014) is commonly used to measure the amount of information lost when an approximate distribution is used to represent a true distribution. A smaller KL divergence indicates that the two distributions are more similar. Thus, in this study, we introduce the SAL based on the KL divergence to align the predicted sentiment distribution with the ground-truth labels, thereby enhancing the accuracy of the emotion intensity predictor. Specifically, we convert $\textbf{I}^{m}$ and ground truth Y into probability distributions $\textbf{P}^{m}$ and $\textbf{P}^{L}$ (Eq.16), then calculate the KL divergence between them (Eq.17). Finally, we scale the KL divergence and incorporate it as an additional regularization term (Eq.18). The scale factor $\beta$ is $10^{3}$. This approach further promotes alignment across different modalities in emotion label prediction.
(16) $\textbf{P}^{m}=\frac{exp(\textbf{I}^{m})}{\sum^{B}_{i=1}exp(\textbf{I}^{m}_{i})},\textbf{P}^{L}=\frac{exp(\textbf{Y})}{\sum^{B}_{i=1}exp(y_{i})},$
(17) $\mathrm{KL}(\textbf{P}^{L}||\textbf{P}^{m})=\frac{1}{B}\sum^{B}_{i=1}\textbf{P}^{L}_{i}(log\textbf{P}^{L}_{i}-log\textbf{P}^{m}_{i}),$
(18) $\mathcal{L}_{SAL}=\sum_{m\in{t,v,a}}\mathrm{KL}(\textbf{P}^{L}||\textbf{P}^{m})\cdot\beta.$
In summary, the total loss $\mathcal{L}_{total}$ for model optimization is expressed as a joint loss defined as follows:
(19) $\mathcal{L}_{total}=\mathcal{L}_{task}+\xi_{1}\cdot\mathcal{L}_{mixMSE}+\xi_{2}\cdot\mathcal{L}_{SAL}.$
where $\xi_{1}$ and $\xi_{2}$ are the weight hyperparameters used to regulate mixed MSE loss and SAL loss, respectively, with the effects of $\xi_{1}$ and $\xi_{2}$ to be evaluated in our ablation study. The SAL can serve as an auxiliary regularization term that not only facilitates the training of the modality-specific predictor $f_{\phi}^{m}$, but also promotes consistency across different modalities.
4. Experiments
To evaluate the proposed approach, we conducted comprehensive experiments on three public benchmark datasets for MSA: CMU-MOSI (MOSI) (Zadeh et al., 2016), CMU-MOSEI (MOSEI) (Zadeh et al., 2018b), and CH-SIMS (SIMS) (Yu et al., 2020). We compare our method against three representative Mixup-based methods (Verma et al., 2019; Venkataramanan et al., 2023; Georgiou et al., 2024) on six state-of-the-art backbones (Zadeh et al., 2017; Liu and Shen, 2018; Tsai et al., 2019; Hazarika et al., 2020; Zhang et al., 2023; Zhuang et al., 2024).
4.1. Datasets
These datasets differ in scale, data collection methods, and linguistic characteristics, thereby offering diverse and challenging testbeds for evaluating model robustness and generalization. We briefly describe the key characteristics of these datasets below:
MOSI
The MOSI (Zadeh et al., 2016) dataset is an English benchmark for emotion recognition, consisting of 2,199 opinion segments from 93 YouTube movie reviews. It offers aligned text (transcripts), audio, and visual (facial) data, annotated for sentiment intensity on a continuous scale from -3 (strongly negative) to +3 (strongly positive) by five independent annotators, with final labels calculated from the average scores. The dataset reflects real-world complexity, including modality incongruity (e.g., sarcasm), and contains about 2.2 hours of video from 89 speakers. Widely utilized in multimodal fusion research, this dataset encompasses pre-extracted features (e.g., acoustic data, facial action units) alongside speaker metadata.
MOSEI
The MOSEI (Zadeh et al., 2018b) dataset includes 23,453 video segments from more than 1,000 different speakers covering over 250 topics. It is the largest multimodal benchmark for sentiment and emotion analysis. The dataset provides aligned text transcripts, audio signals, and visual frames, all annotated by trained annotators with continuous sentiment intensity scores from -3 to 3 like MOSI, and six emotion intensities (happiness, sadness, anger, fear, disgust, surprise). It captures real-world communication nuances, including cross-modal conflicts, and supports research in multimodal fusion, emotion recognition, and sarcasm detection. Preprocessed features of MOSI and MOSEI are available through the CMU-Multimodal SDK (Zadeh et al., 2018c).
SIMS
The SIMS (Yu et al., 2020) dataset is a Chinese multimodal benchmark featuring 2,281 video segments from 200 diverse online videos (vlogs, reviews, dialogues), designed to address real-world modality asynchrony. It provides aligned text transcripts, audio, and visual data with annotations of unified sentiment scores on a continuous scale from -1 (strongly negative) to +1 (strongly positive). As the first Chinese dataset emphasizing cross-modal dynamics, SIMS supports research in asynchronous multimodal fusion, cross-modal interaction modeling, and Mandarin sentiment analysis, with raw videos and preprocessed features publicly available.
4.2. Backbones and Baseline Methods
To ensure a comprehensive comparison, we evaluated the proposed method against three mixup-based methods: Manifold Mixup (Verma et al., 2019), MultiMix (Venkataramanan et al., 2023), and $\mathcal{P}ow$Mix (Georgiou et al., 2024) on six MSA architectures: TFN (Zadeh et al., 2017), LMF (Liu and Shen, 2018), MuIT (Tsai et al., 2019), MISA (Hazarika et al., 2020), ALMT (Zhang et al., 2023), and GLoMo (Zhuang et al., 2024). These architectures represent a diverse spectrum of feature extraction architectures and fusion strategies, covering both early and recent advances in the field, as summarized in Table 1.
Table 1. Description of the six MSA backbones used in the paper
Model TFN (Zadeh et al., 2017) LMF (Liu and Shen, 2018) MuIT (Tsai et al., 2019) MISA (Hazarika et al., 2020) ALMT (Zhang et al., 2023) GLoMo (Zhuang et al., 2024)
Main encoder LSTM LSTM CNN LSTM Transformer Transformer
Fusion method Tensor fusion Low-rank Fusion Cross-attention Self-attention Cross-attention Self-attention
Published In, Year EMNLP, 2017 ACL, 2018 ACL, 2019 ACM MM, 2020 EMNLP, 2023 ACM MM, 2024
4.3. Experimental Settings
Evaluation Metrics
We adopt the evaluation metrics established in the M-SENA framework (Mao et al., 2022) unified framework to ensure consistent and comparable results. These include binary accuracy ($ACC_{2}$) and F1-score, multiclass accuracy ($ACC_{5}$, $ACC_{7}$) and Mean Absolute Error (MAE). Among these, MAE is a regression metric. Classification accuracy values are derived by mapping the regression scores to discrete sentiment categories. The prefix "w-" denotes without neutral-labeled samples. Furthermore, to facilitate intuitive and comprehensive comparison across metrics, we have separately presented the average values (Avg.) of all the classification metrics (ACC and F1-score).
Experimental Details
All experiments were conducted on a high-performance server equipped with an NVIDIA GeForce RTX 3090 Ti GPU. The code was implemented in Python 3.10.18 using the PyTorch 2.7.1 framework with CUDA 12.6 support. For a fair comparison, all models were implemented using publicly available code, and their original experimental settings were retained. Most of the models were evaluated within the M-SENA framework (Mao et al., 2022), ensuring that all the comparison methods are conducted under consistent experimental conditions. In MS-Mix, the base mixing ratio $\lambda_{base}$ is sampled from a symmetric Beta distribution with $\alpha=2.0$ and the similarity threshold $\delta$ is fixed at 0.2. The loss weights $\xi_{1}$ and $\xi_{2}$ are set to 0.7 and 0.5, respectively. Additionally, the number of attention heads $h$ = 4 for the MOSI and MOSEI datasets, and $h$ = 6 for the SIMS dataset.
Table 2. Results of Various Approaches on The MOSI Dataset. Bold: Best performance. Underline: Second-best performance. $\dagger$}: result reported in (Mao et al., 2022). *: result reported in (Georgiou et al., 2024). "-": result was not reported in the original paper.
MODEL w-ACC2 (%) ACC2(%) w-F1-score(%) F1-score(%) ACC5(%) ACC7(%) MAE $\downarrow$ Avg.
TFN 79.08 77.99 79.11 77.95 39.39 34.46 0.947 64.66
+ Manifold Mix 79.27 78.13 79.25 78.06 40.96 35.71 0.927 65.23
+ MultiMix 79.57 78.43 79.60 78.40 39.07 34.69 0.922 64.96
+ $\mathcal{P}ow$Mix 79.27 78.13 77.96 79.15 41.11 36.73 0.919 65.39
+ MS-Mix (ours) 80.18 78.72 80.05 78.52 41.25 34.84 0.904 65.59
LMF 79.18 77.90 79.15 77.80 38.13 33.82 0.950 64.33
+ Manifold Mix 79.42 77.84 79.50 77.86 38.92 35.13 0.948 64.78
+ MultiMix 79.88 78.72 79.96 78.74 38.80 36.32 0.907 65.40
+ $\mathcal{P}ow$Mix 80.49 79.01 80.54 79.00 43.73 38.48 0.895 66.88
+ MS-Mix (ours) 82.16 80.90 82.12 80.81 42.42 36.88 0.893 67.55
MuIT 80.98 79.71 80.95 79.63 42.68 36.91 0.878 66.81
+ Manifold Mix 81.55 79.88 81.48 79.74 43.11 36.55 0.838 67.05
+ MultiMix 81.64 79.01 81.66 79.96 43.38 36.57 0.859 67.04
+ $\mathcal{P}ow$Mix* 81.01 - 80.99 - 40.65 35.00 0.904 -
+ $\mathcal{P}ow$Mix 81.11 79.65 80.64 79.41 41.79 35.25 0.880 66.31
+ MS-Mix (ours) 81.84 79.86 81.80 79.87 43.48 37.51 0.827 67.39
MISA 83.54 81.84 83.58 81.82 47.08 41.37 0.777 69.87
+ Manifold Mix 83.69 82.22 83.69 82.17 46.50 41.11 0.765 69.90
+ MultiMix 83.38 81.34 83.28 81.16 47.38 41.69 0.749 69.71
+ $\mathcal{P}ow$Mix* 83.49 - 83.50 - 48.02 42.65 0.761 -
+ $\mathcal{P}ow$Mix 83.27 81.33 83.41 81.42 47.91 41.87 0.763 69.87
+ MS-Mix (ours) 83.99 82.36 83.96 82.28 48.25 41.98 0.746 70.47
ALMT 84.60 82.65 84.58 82.57 51.75 46.21 0.723 72.06
+ Manifold Mix 83.99 82.22 83.98 82.14 52.62 46.65 0.742 71.93
+ MultiMix 83.69 82.36 83.68 82.30 49.27 44.02 0.751 70.08
+ $\mathcal{P}ow$Mix* 84.30 - 84.30 - 49.24 44.25 0.741 -
+ $\mathcal{P}ow$Mix 84.62 81.63 84.68 81.65 51.96 46.27 0.726 71.80
+ MS-Mix (ours) 84.99 83.22 84.96 82.73 52.19 47.06 0.720 72.36
GLoMo 84.89 83.50 84.92 83.94 53.87 46.13 0.749 72.88
+ Manifold Mix 85.14 83.18 85.07 83.48 53.93 52.27 0.729 73.85
+ MultiMix 84.70 83.42 84.60 83.64 53.82 52.19 0.743 73.73
+ $\mathcal{P}ow$Mix 84.77 84.04 84.71 83.90 53.90 47.53 0.733 73.14
+ MS-Mix (ours) 85.66 83.55 85.58 83.82 54.58 52.90 0.724 74.35
Fig. 3. The t-SNE visualization of the original features and mixed features generated by MS-Mix (a) and $\mathcal{P}$owMix (b) on the MOSEI dataset using the MISA model. We employ a color scheme (blue/red) to differentiate the positive and negative categories, and use transparency to distinguish the original features from the mixed ones. :
4.4. Verification Performance of MS-Mix
This chapter evaluates our proposed method, MS-Mix, on three publicly available MSA datasets: MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018b), and SIMS (Yu et al., 2020). Experiments were conducted using six backbone models, representing both classical (Zadeh et al., 2017; Liu and Shen, 2018; Tsai et al., 2019; Hazarika et al., 2020) and state-of-the-art approaches (Zhang et al., 2023; Zhuang et al., 2024) in terms of model architectures and fusion strategies, to demonstrate the general applicability of MS-Mix across different frameworks. For each model, we compared against classical Mixup variants: Manifold Mixup (Manifold Mix) (Verma et al., 2019) and MultiMix (Venkataramanan et al., 2023), as well as a recently proposed mixup-based method in MSA, $\mathcal{P}ow$Mix (Georgiou et al., 2024). The comparative results are summarized in Tables 2,3,4.
Table 3. Results of Various Approaches on the MOSEI Dataset. Bold: Best performance. Underline: Second-best performance. $\dagger$}: result reported in (Mao et al., 2022). *: result reported in (Georgiou et al., 2024). "-": result was not reported in the original paper.
[…]
TFN 81.89 78.50 81.74 78.96 53.10 51.60 0.573 70.97
+ Manifold Mix 84.31 82.91 84.08 83.03 52.26 51.02 0.572 72.94
+ MultiMix 82.94 80.47 82.80 80.82 53.51 52.39 0.568 72.16
+ $\mathcal{P}ow$Mix 83.46 81.61 83.31 81.87 53.36 51.79 0.562 72.57
+ MS-Mix (ours) 83.63 82.40 83.23 82.36 54.00 52.48 0.557 73.02
LMF 83.48 80.54 83.36 80.94 52.99 51.59 0.576 72.15
+ Manifold Mix 83.74 78.71 83.85 78.98 53.53 52.33 0.563 71.86
+ MultiMix 84.51 81.07 84.44 81.52 54.15 52.65 0.553 73.06
+ $\mathcal{P}ow$Mix 83.38 80.30 83.33 80.79 54.30 53.10 0.559 72.53
+ MS-Mix (ours) 84.65 83.30 84.28 83.26 54.13 52.74 0.554 73.73
MuIT 84.63 81.15 84.52 81.56 54.51 52.84 0.559 73.20
+ Manifold Mix 84.46 81.64 84.54 81.40 54.02 52.51 0.561 73.10
+ MultiMix 84.82 81.31 84.61 81.03 54.79 52.07 0.562 73.11
+ $\mathcal{P}ow$Mix* 84.44 - 84.38 - 54.26 52.75 0.559 -
+ $\mathcal{P}ow$Mix 84.37 81.45 84.56 81.34 54.57 52.45 0.560 73.12
+ MS-Mix (ours) 84.68 81.69 84.64 82.07 54.98 53.09 0.556 73.53
MISA 84.67 80.67 84.66 81.12 53.63 52.05 0.558 72.80
+ Manifold Mix 84.67 80.94 84.67 81.48 53.85 52.05 0.547 72.94
+ MultiMix 84.78 80.49 84.79 81.08 53.68 52.29 0.544 72.85
+ $\mathcal{P}ow$Mix* 84.97 - 84.86 - 54.52 53.00 0.543 -
+ $\mathcal{P}ow$Mix 84.50 81.34 84.75 82.02 53.97 52.87 0.544 73.24
+ MS-Mix (ours) 84.76 83.60 84.95 83.47 54.78 53.06 0.543 74.10
ALMT 85.33 81.13 85.31 81.66 54.71 52.84 0.542 73.50
+ Manifold Mix 86.54 83.69 86.37 83.90 53.38 51.51 0.542 74.23
+ MultiMix 85.64 81.73 85.65 82.26 54.04 54.39 0.531 73.95
+ $\mathcal{P}ow$Mix* 85.85 - 85.94 - 54.89 53.29 0.535 -
+ $\mathcal{P}ow$Mix 85.77 81.25 85.42 81.92 54.17 52.89 0.541 73.57
+ MS-Mix (ours) 85.91 83.22 85.78 83.48 55.18 53.60 0.540 74.53
GLoMo 85.00 83.44 84.88 83.65 54.21 52.47 0.543 73.94
+ Manifold Mix 85.94 83.85 85.80 84.05 55.16 53.09 0.546 74.65
+ MultiMix 84.97 83.90 84.87 83.72 54.55 51.82 0.547 73.97
+ $\mathcal{P}ow$Mix 85.41 83.22 85.29 83.47 55.12 53.18 0.541 74.28
+ MS-Mix (ours) 85.84 84.08 85.72 84.13 55.41 53.27 0.533 74.74
Table 4. Results of Various Approaches on The SIMS Dataset. Bold: Best performance. Underline: Second-best performance. $\dagger$}: result reported in (Mao et al., 2022). *: result reported in (Georgiou et al., 2024). "-": result was not reported in the original paper.
MODEL ACC2 ACC3 ACC5 F1-score MAE $\downarrow$ Avg.
TFN 78.38 65.12 39.30 78.62 0.432 65.36
+ Manifold Mix 79.21 67.83 40.48 78.92 0.443 66.61
+ MultiMix 78.99 66.08 35.45 78.01 0.443 64.63
+ $\mathcal{P}ow$Mix 79.21 67.40 37.42 79.06 0.423 65.77
+ MS-Mix (ours) 81.18 68.71 40.04 81.14 0.421 67.77
LMF 77.77 64.68 40.53 77.88 0.441 65.22
+ Manifold Mix 77.90 65.86 36.98 76.90 0.453 64.41
+ MultiMix 76.59 64.55 36.32 76.70 0.444 63.54
+ $\mathcal{P}ow$Mix 77.86 66.96 39.54 77.33 0.441 65.42
+ MS-Mix (ours) 78.34 65.65 40.61 77.92 0.433 65.63
MuIT 77.46 65.43 34.79 77.57 0.446 63.81
+ Manifold Mix 78.56 65.21 37.86 78.72 0.442 65.09
+ MultiMix 78.24 65.21 38.11 78.41 0.447 64.99
+ $\mathcal{P}ow$Mix* 79.04 - - 78.51 0.437 -
+ $\mathcal{P}ow$Mix 78.64 65.49 37.85 78.90 0.444 65.22
+ MS-Mix (ours) 78.77 65.79 38.73 78.62 0.441 65.48
MISA 76.81 63.24 37.42 76.39 0.467 63.47
+ Manifold Mix 77.19 64.01 36.29 77.05 0.456 63.64
+ MultiMix 77.50 64.18 37.06 77.18 0.437 63.98
+ $\mathcal{P}ow$Mix* 77.35 - - 76.97 0.441 -
+ $\mathcal{P}ow$Mix 77.53 64.03 37.51 77.16 0.439 64.06
+ MS-Mix (ours) 78.54 64.16 37.83 78.18 0.431 64.68
ALMT 76.81 63.02 34.35 76.81 0.449 62.75
+ Manifold Mix 76.37 64.99 36.32 76.66 0.460 63.59
+ MultiMix 77.68 62.63 38.95 78.03 0.433 64.32
+ $\mathcal{P}ow$Mix* 78.91 - - 79.13 0.429 -
+ $\mathcal{P}ow$Mix 78.17 63.06 37.53 78.24 0.431 64.28
+ MS-Mix (ours) 78.34 64.46 37.68 78.21 0.426 64.67
GLoMo 79.09 66.31 37.75 79.20 0.427 65.59
+ Manifold Mix 78.82 66.42 38.02 79.06 0.426 65.58
+ MultiMix 78.89 66.38 38.40 79.27 0.424 65.74
+ $\mathcal{P}ow$Mix 79.03 65.43 38.53 79.16 0.434 65.53
+ MS-Mix (ours) 79.28 67.69 38.06 79.52 0.422 66.14
As shown in Tables 2 to 4, MS-Mix achieves top performance, ranking first or second in all metrics compared to baseline and other mixup-based methods. It is particularly noteworthy that MS-Mix consistently outperforms all competing approaches on the comprehensive average (Avg.) metric across every backbone model.
The consistent outperformance of MS-Mix across diverse backbones and datasets stems from its core design principles, which directly address key limitations in existing mixup-based augmentation methods such as Manifold Mixup (Verma et al., 2019), MultiMix (Venkataramanan et al., 2023), and $\mathcal{P}ow$Mix (Georgiou et al., 2024). Unlike these methods, which often rely on random or offline mixing strategies, MS-Mix incorporates three interconnected mechanisms tailored for MSA: First, the SASS strategy prevents the blending of semantically contradictory samples (e.g., happy and angry) by filtering pairs based on emotional similarity in the latent space. In contrast, the comparative methods perform mixing between random sample pairs regardless of emotional coherence, often resulting in ambiguous or noisy samples that mislead training. By ensuring semantic and label consistency, SASS significantly enhances the quality of augmented data. Second, the SIG mixing module adaptively determines modality-specific mixing ratios based on emotional salience, rather than using random or fixed values as in Manifold Mixup (Verma et al., 2019), MultiMix (Venkataramanan et al., 2023), or leveraging modality importance without emotional context like $\mathcal{P}ow$Mix. This allows MS-Mix to dynamically weight the contribution of each modality, leading to more discriminative mixed features and improved robustness of learned cross-modal representations. Third, the SAL serves as a regularization term that aligns predicted sentiment distributions with ground-truth labels, enhancing prediction consistency across modalities and providing additional supervisory signal. This component is unique to MS-Mix and offers a critical advantage over comparative methods, which lack explicit constraints on emotion-level distribution alignment. As a result, SAL not only stabilizes training but also strengthens the model’s resistance to overfitting.
To further demonstrate the effectiveness of MS-Mix, as shown in Fig. 3, we employed t-SNE (Maaten and Hinton, 2008) to visualize the textual modal feature distributions of both original and mixed features on the MOSEI dataset, using the MISA model as the backbone. The t-SNE visualization demonstrates that features generated by MS-Mix exhibit significantly clearer decision boundaries and more distinct cluster separation compared to those produced by $\mathcal{P}$owMix. This improvement indicates that MS-Mix not only preserves the semantic integrity of the original feature space but also generates more discriminative intermediate representations.
Overall, MS-Mix effectively addresses the semantic confusion and label mismatch issues that hindered previous methods. By integrating three core strategies, our method generates discriminative features with corresponding labels, significantly enhancing the overall quality of augmented data.
4.5. Ablation and Hyperparameter Sensitivity Analysis
To evaluate the impact of key hyperparameters and individual algorithmic components of MS-Mix, we performed ablation studies on three core components using the LMF (Liu and Shen, 2018) model on the MOSI (Zadeh et al., 2016) and SIMS (Yu et al., 2020) datasets. Additionally, hyperparameter sensitivity analyses on the MOSEI dataset (Zadeh et al., 2018b), using the MISA (Hazarika et al., 2020) and GLoMo (Zhuang et al., 2024) models.
In our ablation study, we integrated Manifold Mixup (Verma et al., 2019) into the classic LMF (Liu and Shen, 2018) backbone as the baseline to evaluate the contribution of each proposed component: SASS, SIG, and SAL. Each module was selectively enabled/disabled during training to assess its individual impact on the performance of MS-Mix. Note that since the computation of SAL relies on the output of the SIG module, it cannot be used independently. Results presented in Table 5 show that each of the three components contributes noticeably to performance gains over the baseline, underscoring the importance of every module. Furthermore, by incrementally incorporating these components, we observed progressive improvements in performance, demonstrating not only the individual efficacy of each module but also the significance of synergistic work.
Table 5. Results of the ablation experiments using the LMF method on the MOSI database. Bold: Best performance.
[…]
79.42 77.84 79.50 77.86 38.92 35.13 0.948 64.78
79.98 78.41 80.08 78.60 38.86 35.26 0.921 65.20
80.94 78.91 80.93 79.00 40.48 35.25 0.925 65.92
81.48 79.71 81.58 79.38 41.54 35.37 0.949 66.51
81.41 78.03 81.41 79.61 40.48 36.00 0.919 66.16
82.16 80.90 82.12 80.81 42.42 36.88 0.893 67.55
Fig. 4. The impact of different parameter values. (a) Accuracy under different similarity thresholds $\delta$. (b) Accuracy under different combinations of $\xi_{1}$ and $\xi_{2}$. (c) Accuracy under different $\alpha$. (d) Accuracy under different attention heads $h$. :
To comprehensively analyze the sensitivity of key hyperparameters, we conducted extensive experiments using both the top-performing GLoMo model (Zhuang et al., 2024) and the classical MISA architecture (Hazarika et al., 2020) on the largest English dataset MOSEI (Zadeh et al., 2018b) and the Chinese benchmark SIMS (Yu et al., 2020). As illustrated in Fig. 4, our evaluation reveals how critical hyperparameters affect model performance. Using the MISA model on MOSEI, we first examined the similarity selection threshold $\delta$ (Fig. 4(a)). The results demonstrate that at $\delta=0.2$, the method maintains an optimal balance, incorporating sufficient samples while effectively filtering out emotionally contradictory ones, thus achieving peak performance. Notably, even under suboptimal $\delta$ configurations, MS-Mix consistently surpasses Manifold Mixup (Verma et al., 2019). We further evaluated the loss weights $\xi_{1}$ and $\xi_{2}$ with the GLoMo model on MOSEI (Fig. 4(b)). The experimental outcomes confirm the rationality of our weight selection, showing that $\xi_{1}=0.7$ and $\xi_{2}=0.5$ yield the best performance. Additionally, we assessed the shape parameter $\alpha$ of the Beta distribution using MISA on MOSEI (Fig. 4(c)), finding $\alpha=2.0$ to be optimal. The influence of the number of attention heads $h$ was also investigated across datasets (Fig. 4(d)): $h=4$ delivers the best results on MOSEI, while $h=6$ is most effective on SIMS. These findings collectively validate the rationale behind our parameter configurations.
4.6. Occlusion Experiment
In the random occlusion experiment, we compared the performance of our method against baseline approaches (Verma et al., 2019; Venkataramanan et al., 2023; Georgiou et al., 2024) under both clean and noisy data conditions across MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018b), and SIMS (Yu et al., 2020). Specifically, we randomly masked out 0% to 40% of the input data during training of the LMF (Liu and Shen, 2018) model to investigate the effect of mixup-based augmentation on model robustness.
Fig. 5. The performance of different mixup-based methods at different occlusion ratios on three datasets. :
As illustrated in Fig. 5, the experimental results demonstrate that the introduction of noise influences model performance across all datasets. A noteworthy phenomenon emerges in which moderate levels of random masking (e.g., 10% or 20%) unexpectedly improve performance, rather than degrade it. Beyond these optimal masking ratios, however, model performance deteriorates. Despite these variations, MS-Mix consistently achieves the highest average accuracy across all evaluated conditions, demonstrating its advantages in optimizing the model for handling noisy data.
5. CONCLUSION
In this paper, we propose MS-Mix, a novel and adaptive data augmentation framework specifically designed for MSA. Unlike traditional mixup-based methods, which often rely on random and offline mixing strategies, MS-Mix introduces three key emotion-aware mechanisms to enhance the semantic consistency of augmented samples. First, the SASS module ensures that only emotionally consistent samples are mixed, effectively avoiding mixing feature pairs with contradictory emotions. Second, the SIG mixing module adaptively determines mixing ratios of each modality based on emotional salience, promoting the generation of samples with clearer classification boundaries. Third, the SAL aligns the predicted emotion distributions with ground-truth labels through the KL-divergence minimization, serving as an effective regularizer to improve generalization. Empirical evaluation of three benchmark datasets and six diverse backbone architectures has shown that MS-Mix achieves state-of-the-art performance.
In the future, we will consider extending MS-Mix to other multimodal tasks, such as mental health monitoring and human-computer interaction, which could leverage its ability to maintain semantic consistency across a wider range of applications. We also intend to integrate self-supervised or semi-supervised learning strategies into MS-Mix to alleviate the reliance on large annotated datasets and enhance its applicability in low-resource settings. Finally, from an efficiency perspective, we will focus on developing lightweight variants of the sample selection and fusion mechanisms to reduce computational costs and support deployment on resource-constrained devices.
6. Acknowledgments
This work was supported by the Science and Technology Innovation Key R&D Program of Chongqing (No. CSTB2024TIAD-STX0023, CSTB2023TIAD-STX0015, and CSTB2023TIAD-STX0031), and in part by the Natural Science Foundation of China (No. 62372427).

### v2

@@
-MS-Mix: Unveiling the Power of Mixup for Multimodal Sentiment Analysis
-Abstract.
-Multimodal Sentiment Analysis (MSA) aims to identify and interpret human emotions by integrating information from heterogeneous data sources such as text, video, and audio.
-While deep learning models have advanced in network architecture design, they remain heavily limited by scarce multimodal annotated data.
-Although Mixup-based augmentation improves generalization in unimodal tasks, its direct application to MSA introduces critical challenges: random mixing often amplifies label ambiguity and semantic inconsistency due to the lack of emotion-aware mixing mechanisms.
+MS-Mix: An Adaptive Emotion-Sensitive Mixup Framework for Multimodal Sentiment Analysis
+Abstract
+Multimodal Sentiment Analysis (MSA) integrates complementary features from text, video, and audio to achieve robust emotion understanding.
+Yet, models for this task suffer from severe data scarcity and high annotation costs.
+This severely restricts their practical engineering applications.
+Although Mixup-based augmentation improves generalization in unimodal settings, its naive application to MSA often produces semantically inconsistent mixtures and amplified label noise due to random sample mixing and fixed mixing ratios that ignore emotional semantics.
 To overcome these issues, we propose MS-Mix, an adaptive, emotion-sensitive augmentation framework that automatically optimizes sample mixing in multimodal settings.
-The key components of MS-Mix include: (1) a Sentiment-Aware Sample Selection (SASS) strategy that effectively prevents semantic confusion caused by mixing samples with contradictory emotions.
-(2) a Sentiment Intensity Guided (SIG) module using multi-head self-attention to compute modality-specific mixing ratios dynamically based on their respective emotional intensities.
-(3) a Sentiment Alignment Loss (SAL) that aligns the prediction distributions across modalities, and incorporates the Kullback-Leibler-based loss as an additional regularization term to train the emotion intensity predictor and the backbone network jointly.
-Extensive experiments on three benchmark datasets with six state-of-the-art backbones confirm that MS-Mix consistently outperforms existing methods, establishing a new standard for robust multimodal sentiment augmentation.
-The source code is available at: https://github.com/HongyuZhu-s/MS-Mix.
-Multimodal sentiment analysis, Data augmentation, Regularization
-copyright: acmlicensedcopyright: acmlicensedjournal: TOMMjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: XXXXXXX.XXXXXXXccs: Computing methodologies Neural networksccs: Computing methodologies Regularization
-1. Introduction
-The perception and understanding of human emotions by Artificial Intelligence (AI) is of great significance for technologies such as human-computer interaction (Hibbeln et al., 2017) and multimedia computing(Li et al., 2019).
-Multimodal Sentiment Analysis (MSA) has emerged as a critical research frontier in this endeavor (Dumpala et al., 2019; Liu et al., 2025; Mao et al., 2022; Zhang et al., 2023; Zhuang et al., 2024; Xiao et al., 2025).
-MSA aims to integrate and interpret complementary emotional cues from textual, acoustic, and visual modalities to achieve a more robust and accurate understanding of sentiment (Manning et al., 2014).
-The significance of MSA is underscored by its wide-ranging applications, including but not limited to opinion mining on social media (Chandrasekaran et al., 2021), empathetic chatbot design (Yang et al., 2019), and automated mental health assessment (Singh et al., 2024).
-By moving beyond unimodal analysis, MSA offers the potential to capture the complex and often incongruent nature of human affective expression, thereby providing a more holistic view of sentiment.
-Benefiting from the rapid advances in Deep Learning (DL), the field of MSA has witnessed substantial progress in the development of neural architectures.
-Techniques such as cross-modal transformers (Tsai et al., 2019) and advanced fusion mechanisms (Zhuang et al., 2024) have demonstrated remarkable capability in modeling inter-modal interactions and extracting discriminative features.
-Nevertheless, despite these architectural advancements, the performance of data-driven DL models remains fundamentally constrained by the scale and quality of annotated data (Zhang et al., 2018).
-This data scarcity often results in overfitting and limited generalization ability (Srivastava et al., 2014).
-Furthermore, the construction of MSA datasets is costly, as it necessitates the collection of reliable human-annotated multimodal sentiment dataset.
-Fig.
-1.
-The differences between MS-Mix and traditional mixup methods represented by Manifold Mixup (Zhang et al., 2018).
+The key components of MS-Mix include: (1) a Sentiment-Aware Sample Selection strategy that filters incompatible pairs by measuring semantic similarity in the latent space, preventing mixtures of contradictory emotions.
+(2) a Sentiment Intensity Guided module using multi-head self-attention, that dynamically computes modality-specific mixing ratios conditioned on emotional salience.
+(3) a Sentiment Alignment Loss based on Kullback-Leibler divergence that regularizes the model by aligning predicted sentiment distributions across modalities with ground-truth labels, improving discrimination and consistency.
+Extensive experiments on three benchmark datasets with six state-of-the-art backbones confirm that MS-Mix consistently outperforms existing methods, thereby improving the robustness and practical applicability of MSA in engineering contexts.
+The source code is available at an anonymous link: https://anonymous.4open.science/r/MS-Mix-review-0C72.
+keywords:
+Affective computing , Multimodal sentiment analysis ,
+Data augmentation , Regularization
+journal: Nuclear Physics B
+[…]
+[1]organization=Chongqing Institute of Green Intelligent Technology, Chinese Academy of Sciences; Chongqing School, University of Chinese Academy of Sciences, addressline=400714, city=Chongqing, country=China
+[…]
+[2]organization=School of Engineering, Westlake University, city=Zhejiang, country=China
+[…]
+[3]organization=Telecom SudParis, Institute Polytechnique de Paris, city=Palaiseau, country=France
+1 Introduction
+The perception and understanding of human emotions by Artificial Intelligence (AI) are of great significance for technologies such as human-computer interaction [1], and multimedia computing [2].
+Multimodal Sentiment Analysis (MSA) has emerged as a critical research frontier in this endeavor [3, 4, 5, 6, 7].
+MSA aims to integrate and interpret complementary emotional features from textual, acoustic, and visual modalities to achieve a more robust and accurate understanding of sentiment [8].
+The significance of MSA is underscored by its wide-ranging applications, including but not limited to opinion mining on social media [9], empathetic chatbot design [10], and mental health assessment [11].
+By moving beyond unimodal analysis, MSA can capture the complex, often incongruent nature of human affective expression, thereby providing a more holistic view of sentiment.
+Benefiting from rapid advances in Deep Learning (DL), the field of MSA has witnessed substantial progress in developing neural architectures.
+Techniques such as cross-modal transformers [12] and advanced fusion mechanisms [6] have demonstrated remarkable capability in modeling inter-modal interactions and extracting discriminative features.
+Nevertheless, despite these architectural advancements, the performance of data-driven DL models remains fundamentally constrained by the scale and quality of annotated data [13].
+This data scarcity often results in overfitting and limited generalization ability [14].
+Furthermore, constructing MSA datasets is costly, as it requires the collection of reliable, human-annotated multimodal sentiment datasets.
+Above all, these issues severely restrict the practical engineering applications of MSA.
+Figure 1: The differences between MS-Mix and traditional mixup methods represented by Manifold Mixup [13].
 Left: a.
-The traditional mixup method employs random selection of samples and an offline mixing ratios optimization strategy.
+The traditional methods employ random sampling and an offline optimization of mixing ratios.
 b.
@@
 MS-Mix.
-Right: MS-Mix can generate augmented samples that better align with the original distribution on the MOSEI dataset.
-Where the KL represents the Kullback-Leibler divergence.
-:
-To mitigate the data scarcity issue, the mixup strategy, a data augmentation technique, was introduced to generate virtual training samples via convex interpolations between original data points and their corresponding labels (Zhang et al., 2018).
-However, existing offline methods, including general domain techniques such as CutMix (Yun et al., 2019) and SaliencyMix (Uddin et al., 2021), as well as multimodal specific approaches like $\mathcal{P}ow$Mix (Georgiou et al., 2024) and MultiMix (Venkataramanan et al., 2023), often depend on randomly paired samples and offline optimization strategies, thereby neglecting the underlying emotional semantics.
+Right: MS-Mix can generate augmented samples that better align with the original label distribution.
+To mitigate data scarcity, the mixup strategy, a data augmentation technique, was introduced to generate virtual training samples by convexly interpolating between original data points and their corresponding labels [13].
+However, existing offline methods, including general domain techniques such as CutMix [15] and SaliencyMix [16], as well as multimodal specific approaches like $\mathcal{P}ow$Mix [17] and MultiMix [18], often depend on randomly paired samples and offline optimization strategies, thereby neglecting the underlying emotional semantics.
 This may result in the blending of semantically inconsistent samples, e.g., combining a positive and a negative sample can introduce semantic confusion and label mismatch.
-Moreover, these methods typically employ fixed or uniformly distributed mixing ratios, which fail to adapt to the varied emotional intensities across different modalities.
+Moreover, these methods typically use fixed or uniformly distributed mixing ratios, which do not adapt to the varying emotional intensities across modalities.
 As a consequence, the quality and efficacy of the augmented data are substantially compromised.
-To address these limitations, we propose MS-Mix, a novel and adaptive data augmentation framework specifically designed for multimodal sentiment analysis.
+To address these limitations, we propose MS-Mix, a novel, adaptive data augmentation framework designed for multimodal sentiment analysis.
 As shown in Fig.
-1, we have visualized the distribution of the generated labels.
+1, we have visualized the distribution of the generated labels, and MS-Mix has a lower Kullback- Leibler (KL) divergence.
 It can be observed that the label distribution generated by MS-Mix is closer to the real situation.
@@
 The SAL acts as a regularization term, enhancing the backbone network’s discriminative power and improving the consistency of representations from augmented samples.
-Experimental results conducted on the real-world MSA datasets demonstrate that MS-Mix significantly outperforms existing methods across multiple baselines and backbone architectures.
+Experimental results on real-world MSA datasets show that MS-Mix significantly outperforms existing methods across multiple baselines and backbone architectures.
 The main contributions of this paper are summarized as follows:
-  • We design an SASS strategy that leverages semantic similarity in the latent space to filter out incompatible sample pairs, effectively preventing mixtures of samples with contradictory emotions.
+  • 1.
+    We design a SASS strategy that leverages semantic similarity in the latent space to filter out incompatible sample pairs, effectively preventing mixtures of samples with contradictory emotions.
+  • 2.
     We introduce a SIG mixing module, implemented via a multi-head self-attention mechanism, to dynamically determine modality-specific mixing ratios based on emotional salience.
-  • We propose SAL, a Kullback-Leibler (KL) divergence-based regularization that aligns predicted sentiment distributions with ground-truth labels, thereby enhancing both model discrimination and consistency.
-  • We conduct extensive experiments on three popular MSA datasets and six backbone architectures, demonstrating that MS-Mix significantly outperforms existing methods.
-The remainder of this paper is organized as follows.
-Section 2 reviews related methods.
+  • 3.
+    We propose SAL, a Kullback-Leibler divergence-based regularization that aligns predicted sentiment distributions with ground-truth labels, thereby enhancing both model discrimination and consistency.
+The remainder of this paper is organized as follows: Section 2 reviews related works.
 The proposed MS-Mix method is detailed in Section 3.
-The experimental results are presented in Section 4, and the concluding remarks are provided in Section 5.
-2. Related Work
-2.1. Multimodal Sentiment Analysis
-Sentiment analysis systems are divided into single-modal systems and multi-modal systems.
-Compared with single-modal systems, which analyze the sentiment state through a single data source, MSA systems can utilize complementary information from various modalities such as text, video, and audio, thereby providing an effective method for comprehensive sentiment analysis (Geetha et al., 2024).
+The experimental results and visualizations are presented in Section 4, and the conclusions are provided in Section 5.
+2 Related Work
+2.1 Multimodal Sentiment Analysis
+Sentiment analysis systems are divided into single-modal systems and multimodal systems.
+Compared with single-modal systems, which analyze sentiment using a single data source, MSA systems can leverage complementary information across multiple modalities, such as text, video, and audio, thereby providing a more comprehensive approach to sentiment analysis [19].
 For instance, Cimtay et al.
-(Cimtay et al., 2020) achieved emotion recognition by extracting data features from Galvanic Skin Response (GSR), Electroencephalogram (EEG), and facial expressions.
-Prior research in MSA has predominantly advanced through improving two main paradigms: feature fusion strategy-centric (Zadeh et al., 2017; Liu and Shen, 2018; Zadeh et al., 2018a; Tsai et al., 2019; Lv et al., 2021; Zhuang et al., 2024) and feature encoder method-centric (Liang et al., 2021; Han et al., 2021; Yang et al., 2022; Zhang et al., 2023) approaches.
+[20] achieved emotion recognition by extracting data features from Galvanic Skin Response (GSR), Electroencephalogram (EEG), and facial expressions.
+Prior research in MSA has predominantly focused on improving two main paradigms: feature-fusion strategy-centric [21, 22, 23, 12, 24, 6] and feature-encoder method-centric [25, 26, 27, 5] approaches.
 Fusion strategy-centric approaches in MSA primarily focused on designing effective fusion strategies to integrate features from different modalities.
 Zadeh et al.
-(Zadeh et al., 2017) introduced the Tensor Fusion Network (TFN), which explicitly models inter-modal interactions through the 3-fold cartesian product.
-Building on this work, Z Liu et al.
-(Liu and Shen, 2018) proposed Low-rank Multimodal Fusion (LMF), an efficient approach that enhances computational performance via parallel tensor and weight decomposition.
+[21] introduced the Tensor Fusion Network (TFN), which explicitly models intermodal interactions via the 3-fold Cartesian product.
+Building on this work, Z.
+Liu et al.
+[22] proposed Low-rank Multimodal Fusion (LMF), an efficient approach that improves computational performance through parallel tensor and weight decomposition.
 By leveraging modality-specific low-rank factors, LMF achieves effective multimodal fusion with significantly reduced computational cost.
-MuIT (Tsai et al., 2019) incorporates cross-modal attention, exploring the long-range dependency relationships among cross-modal elements in multimodal datasets, and achieving multimodal fusion on unaligned multimodal datasets.
-Based on the same motivation and algorithm, Huang et al.
-(Huang et al., 2020) employed cross-modal attention to learn the long-range dependencies between the visual and audio modalities, achieving better performance than decision-level and feature-level fusion.
+MuIT [12] incorporates cross-modal attention, explores long-range dependencies among cross-modal elements in multimodal datasets, and achieves multimodal fusion across unaligned datasets.
 A recent advancement in MSA is the attempt to decouple the features into shared and unique information.
 For instance, Li et al.
-(Li et al., 2023) proposed a Decoupled Multimodal Distillation (DMD) method, which is capable of distilling cross-modal knowledge in the decoupled feature spaces and alleviating the problem of inherent multimodal heterogeneity.
-The Global Local Modal (GLoMo) (Zhuang et al., 2024) Fusion framework integrates multiple local representations within each modality and effectively combines local and global information.
-For feature encoder-centric approaches, the primary objective is to enhance the feature representation of each modality.
+[28] proposed a Decoupled Multimodal Distillation (DMD) method that distills cross-modal knowledge in decoupled feature spaces and alleviates the inherent multimodal heterogeneity.
+The Global Local Modal (GLoMo) [6] Fusion framework integrates multiple local representations within each modality and effectively combines local and global information.
+In feature encoder-centric approaches, the primary objective is to improve the feature representation for each modality.
 Yang et al.
-(Yang et al., 2022) conceptualize multimodal representation learning as a form of domain adaptation, employing adversarial learning to model both modality-invariant and modality-specific subspaces within multimodal fusion.
-Han et al.
-(Han et al., 2021) introduced the MMIM framework, which enhances multimodal fusion performance through hierarchical mutual information maximization.
+[26] conceptualize multimodal representation learning as a form of domain adaptation, employing adversarial learning to model both modality-invariant and modality-specific subspaces within multimodal fusion.
 To address the challenge of multimodal heterogeneity in MSA, Li et al.
-(Lin and Hu, 2023) conceptualized the learning process of each modality as a set of sub-tasks and introduced a novel Multi-Task Momentum Distillation (MTMD) method.
+[27] conceptualized the learning process of each modality as a set of sub-tasks and introduced a novel Multi-Task Momentum Distillation (MTMD) method.
 This approach effectively reduces the discrepancies between modalities and enhances representation consistency.
-Furthermore, to suppress sentiment-irrelevant and conflicting information across modalities, Zhang et al.
-(Zhang et al., 2023) proposed the Adaptive Language-guided Multimodal Transformer (ALMT), which learns a unified hypermodality representation guided by language at multiple scales.
-In recent work, Li et al.
-(Li et al., 2025) proposed the MM-PEAR-CoT framework, which enhances the reliability of multimodal sentiment analysis by using cross-modal filtering and fusion to suppress irrational reasoning steps generated by large language models.
-Despite these advancements, MSA systems remain constrained by the scarcity and high annotation cost of multimodal data, often leading to overfitting and limited generalization (Venkataramanan et al., 2023; Georgiou et al., 2024).
+Furthermore, to suppress sentiment-irrelevant and conflicting information across modalities, in recent work, Zhang et al.
+[5] proposed the Adaptive Language-guided Multimodal Transformer (ALMT), which learns a unified hypermodality representation guided by language at multiple scales.
+Despite these advancements, MSA systems remain constrained by the scarcity and high annotation costs of multimodal data, which often lead to overfitting and limited generalization [18, 17].
 This highlights the need for effective data augmentation techniques specifically designed for multimodal emotional data.
-2.2. Mixup-based Augmentation
-To mitigate data scarcity and enhance generalization, data augmentation techniques—particularly Mixup and its variants—have been extensively adopted in DL (Kim et al., 2020a; Liu et al., 2022; Qin et al., 2024a, b; Jin et al., 2024).
-The original Mixup method (Zhang et al., 2018) operates in the input space by generating virtual samples through linear interpolation between two randomly selected data points and their corresponding labels.
+2.2 Mixup-based Augmentation
+To mitigate data scarcity and improve generalisation, data augmentation techniques, particularly Mixup and its variants, have been widely adopted in DL [29, 30, 31, 32, 33].
+The original Mixup method [13] operates in the input space by generating virtual samples through linear interpolation between two randomly selected data points and their corresponding labels.
 This simple yet effective strategy encourages models to behave linearly across classes and improves robustness.
-Subsequent extensions such as Manifold Mixup (Verma et al., 2019) generalize the interpolation operation to hidden representations, further enhancing the smoothness of decision boundaries.
-Similarly, in the latent space, the Mixup-Transformer (Sun et al., 2020) encodes a sentence using a Transformer and then linearly interpolates the representations to generate mixed samples for classification.
-In recent state-of-the-art approaches (Liu et al., 2022; Qin et al., 2024a, b), mixing strategies are no longer manually designed based on prior knowledge or saliency information but instead adaptively generate mixed samples using learnable mixing ratios and feature representations in an end-to-end manner.
+Subsequent extensions, such as Manifold Mixup [34], generalize the interpolation operation to hidden representations, further enhancing the smoothness of decision boundaries.
+Similarly, in the latent space, the Mixup-Transformer [35] encodes a sentence using a Transformer and then linearly interpolates the representations to generate mixed samples for classification.
+In the latest work, AdvMixUp [33] introduces a sample-dependent, feature-level interpolation mask generated by a compact Mask Generator, optimized via adversarial training (min-max framework) to produce challenging ”hard” mixed samples near decision boundaries.
+Recent state-of-the-art approaches [30, 31, 33] that use mixing strategies no longer rely on manually designed rules based on prior knowledge or saliency information, but instead adaptively generate mixed samples using learnable mixing ratios and feature representations in an end-to-end manner.
 In the multimodal domain, several works have extended Mixup to leverage cross-modal interactions.
-For instance, MultiMix (Venkataramanan et al., 2023) performs independent Mixup operations within each modality and integrates the augmented representations through late fusion.
-VLMixer (Wang et al., 2022) integrates cross-modal CutMix (Yun et al., 2019) with contrastive learning to convert uni-modal text inputs into multi-modal representations of text and image, thereby improving instance-level alignment across modalities.
+For instance, MultiMix [18] performs independent Mixup operations within each modality and integrates the augmented representations through late fusion.
+VLMixer [36] integrates cross-modal CutMix [15] with contrastive learning to convert uni-modal text inputs into multi-modal representations of text and image, thereby improving instance-level alignment across modalities.
 To enhance robustness against missing modalities, Lin et al.
-(Lin and Hu, 2024) proposed Multi-Modal Mixup ($M^{3}$ixup), which extends mixup to both representation and contrastive loss levels for improved cross-modal alignment and dynamics capture.
-More recently, as an improvement of MultiMix, $\mathcal{P}$owMix (Georgiou et al., 2024) introduced a weight-aware mixing strategy that dynamically modulates mixing coefficients based on the estimated importance of each modality, enabling more flexible and effective multimodal augmentation.
-However, most existing methods rely on random sample pairing and offline mixing schemes, which overlook the semantic structure and emotional coherence of multimodal data.
+[37] proposed Multi-Modal Mixup ($M^{3}$ixup), which extends mixup to both representation and contrastive loss levels to improve cross-modal alignment and capture of dynamics.
+More recently, as an improvement of MultiMix, $\mathcal{P}$owMix [17] introduced a weight-aware mixing strategy that dynamically modulates mixing coefficients based on the estimated importance of each modality, enabling more flexible and effective multimodal augmentation.
+However, most existing methods rely on random sampling and offline mixing, which overlook the semantic structure and emotional coherence of multimodal data.
 This often results in semantically inconsistent mixtures, such as mixing samples with opposing emotions, which introduces label noise and hinders model performance.
@@
 Therefore, we propose MS-Mix, a novel mixup method that effectively avoids the mixing of contradictory emotions and adaptively generates discriminative multimodal emotional features.
-Fig.
-2.
-An overview of the overall structure of the proposed MS-Mix framework.
+Figure 2: An overview of the proposed MS-Mix framework.
 (a) The SASS strategy computes the emotional semantic distance between samples.
@@
 These components work collaboratively to enhance multimodal representation learning.
-:
-3. Methodology
-In this section, we first introduce the task definition and the overview of our MS-Mix framework. We then describe the details of each module in MS-Mix.
-3.1. Task Definition and Framework Overview
-Given a MSA dataset with $n$ samples, $\textbf{X}=(\textbf{x}_{1},\textbf{x}_{2},...,\textbf{x}_{n})$} and $\textbf{Y}=(y_{1},y_{2},...,y_{n})$} are the data and labels.
-For the $i_{th}$ sample $\textbf{x}_{i}\in\textbf{X}$, where $\textbf{x}_{i}$ comprises data from the text (t), video (v), and audio (a) modalities, denoted as $\{\textbf{x}^{t}_{i},\textbf{x}^{v}_{i},\textbf{x}^{a}_{i}\}$.
-The goal of the MSA task is to learn a mapping function $\mathcal{F}:\textbf{X}\mapsto\textbf{Y}$ to predict the occurrence of each emotion category.
-MS-Mix is a novel data augmentation method in the latent space, acting on the multimodal features output by the encoder, and is applied before feature fusion.
+3 Methodology
+In this section, we first introduce the task definition and the overview of our MS-Mix framework.
+We then describe the details of each module in MS-Mix.
+The complete MS-Mix process is presented in Algorithm 1.
+Input: Data X and label Y, backbone $f_{\theta}$, prediction head $f_{\psi}$,
+hyperparameters $\alpha$, $\delta$, $\beta$
+1
+2Initialize parameters;
+3
+4for each epoch do
+5   Sample mini-batch $B$;
+6
+7 // Step 1: Sentiment-Aware Sample Selection Extract features $\textbf{Z}^{m}=f_{\theta}(\textbf{X})$ for each modality;
+8   Compute averaged similarity across modalities;
+9   Select consistent pairs where $s_{ij}>\delta$;
+10
+11 for each consistent pair $((x_{i},y_{i}),(x_{j},y_{j}))$ do
+12    // Step 2: Intensity-Guided Mixing
+13      Predict intensities $I^{m}$ via $\textbf{Z}^{m}$;
+14      Obtain modality specific mixing ratios $\lambda^{m}_{i,j}$, and mixing ratios $\lambda^{L}_{i,j}$ of labels;
+15
+16     Generate mixed features $\hat{\textbf{z}}^{m}=\lambda^{m}_{i,j}\textbf{z}_{i}^{m}+(1-\lambda^{m}_{i,j})\textbf{z}_{j}^{m}$, and mixed label $\hat{y}=\lambda^{L}_{i,j}y_{i}+(1-\lambda^{L}_{i,j})y_{j}$ ;
+17
+18    // Step 3. Losses & Update
+[…]
+[…]
+21      Alignment loss: $\mathcal{L}_{\text{SAL}}\leftarrow\text{KL}(P^{L}||P^{m})$;
+22      Total loss: $\mathcal{L}_{total}=\mathcal{L}_{task}+\xi_{1}\cdot\mathcal{L}_{mixMSE}+\xi_{2}\cdot\mathcal{L}_{SAL}$;
+23
+24     Update $\theta$ by minimizing $\mathcal{L}_{total}$;
+25
+26   end for
+27
+28 end for
+29return Trained $f_{\psi}\circ f_{\theta}$;
+Algorithm 1 Training Procedure of MS-Mix
+3.1 Task Definition and Framework Overview
+Given a MSA dataset with $n$ samples, $\textbf{X}=(\textbf{x}_{1},\textbf{x}_{2},...,\textbf{x}_{n})$} and $\textbf{Y}=(y_{1},y_{2},...,y_{n})$} where $y_{i}$ are continuous sentiment intensity scores, each $\mathbf{x}_{i}\in\mathbf{X}$ comprises data from the text ($t$), video ($v$), and audio ($a$) modalities, denoted as $\{\mathbf{x}_{i}^{t},\mathbf{x}_{i}^{v},\mathbf{x}_{i}^{a}\}$.
+The goal of the MSA task is to learn a mapping function $\mathcal{F}:\mathbf{X}\mapsto\mathbf{Y}$ to predict the continuous intensity of each emotion category.
+This mapping is realized by a backbone $f_{\theta}$ followed by a prediction head $f_{\psi}$.
+To address data scarcity and enhance generalization in this regression-based MSA task, we propose MS-Mix.
+This novel latent-space data augmentation method operates on the multimodal features output by the encoder and is applied before feature fusion.
 As illustrated in Fig.2, the input to MS-Mix consists of multimodal features derived from the model’s encoder outputs.
@@
 Additionally, the SAL serves as an auxiliary regularization term to jointly optimize the entire network and the SIG module.
-As a common starting point for many mixup-based augmentation methods, vanilla mixup (Zhang et al., 2018) employs a ratio $\lambda$ to construct a mixed sample ($\hat{\textbf{x}},\hat{y}$) by performing global linear interpolation directly on the sample pair $\{(\textbf{x}_{i},y_{i}),(\textbf{x}_{j},y_{j})\}$:
-(1) $\hat{\textbf{x}}=\lambda\cdot\textbf{x}_{i}+(1-\lambda)\cdot\textbf{x}_{j},$
+As a common starting point for many mixup-based augmentation methods, vanilla mixup [13] employs a ratio $\lambda$ to construct a mixed sample ($\hat{\textbf{x}},\hat{y}$) by performing global linear interpolation directly on the sample pair $\{(\textbf{x}_{i},y_{i}),(\textbf{x}_{j},y_{j})\}$:
+$\hat{\textbf{x}}=\lambda\cdot\textbf{x}_{i}+(1-\lambda)\cdot\textbf{x}_{j},$ (1)
 $\hat{y}=\lambda\cdot y_{i}+(1-\lambda)\cdot y_{j},$
@@
 We define the process of Eq.1 as follows: Given the sample mixup function $\mathrm{H}(\cdot)$, the label mixup function $\mathrm{G}(\cdot)$, and a mixing ratio $\lambda$, we can generate the mixed sample ($\hat{\textbf{x}},\hat{y}$) with $\hat{\textbf{x}}=\mathrm{H}(\textbf{x}_{i},\textbf{x}_{j},\lambda)$, $\hat{y}=\mathrm{G}(y_{i},y_{j},\lambda)$.
-Mixing of original data can easily lead to sentiment semantic confusion in generated data (Verma et al., 2019).
+Mixing of original data can easily lead to sentiment semantic confusion in generated data [34].
 Therefore, for each batch of data, we choose to learn different mixing ratios $\lambda^{m}$ to mix the features $\textbf{Z}^{m}=f_{\theta}(\textbf{X})\in\mathbb{R}^{B\times d^{m}}$ of each modal in the latent space separately (Eq.
 2), and calculate the label ratio $\lambda^{L}$ to mix the labels as follows:
-(2) $\hat{\textbf{z}}^{m}=\mathrm{H}(f_{\theta}(\textbf{x}_{i}^{m}),f_{\theta}(\textbf{x}_{j}^{m}),\lambda^{m}),\textbf{x}^{m}\in\textbf{X},$
-(3) $\hat{y}=\mathrm{G}((y_{i},y_{j}),\bm{\lambda}^{L}),y\in\textbf{Y}.$
-where $\hat{\textbf{z}}^{m}$ represents the the mixed features, $m\in\{t,v,a\}$ represents a modality belonging to the set of the three modalities in the dataset, $B$ is the mini-batch size, $d^{m}$ is the hidden space dimension of modality $m$, and $f_{\theta}$ is the backbone.
-Finally, the mixed features and the original features will be fused into a whole set for the subsequent feature fusion process.
-3.2. Sentiment-Aware Sample Selection
+$\hat{\textbf{z}}^{m}=\mathrm{H}(f_{\theta}(\textbf{x}_{i}^{m}),f_{\theta}(\textbf{x}_{j}^{m}),\lambda^{m}),\textbf{x}^{m}\in\textbf{X},$ (2)
+$\hat{y}=\mathrm{G}((y_{i},y_{j}),\bm{\lambda}^{L}),y\in\textbf{Y},$
+where $\hat{\textbf{z}}^{m}$ represents the mixed features, $m\in\{t,v,a\}$ represents the set of the three modalities in the dataset, $B$ is the mini-batch size, $d^{m}$ is the hidden space dimension of modality $m$, and $f_{\theta}$ is the backbone.
+Finally, the mixed and original features will be fused into a single set for subsequent feature fusion.
+3.2 Sentiment-Aware Sample Selection
 Emotional expressions exhibit significant variation across samples.
-While random in-batch mixing enhances the feature smoothness, it often ignores semantic consistency in the emotional space.
+While random in-batch mixing enhances feature smoothness, it often fails to preserve semantic consistency in the emotional space.
 This may lead to the mixing of two semantically opposite samples with strong emotion into a mixed sample with a neutral label and semantic confusion features.
-Therefore, we proposed the SASS strategy based on the emotional semantic distance, and screened the feature samples within a batch before mixing.
-Specifically, we use the feature cosine distance between each sample to represent the semantic information similarity (Pennington et al., 2014), and mix the sample pairs with a similarity greater than the threshold $\delta$ within each batch.
-This ensures that only samples with analogous emotions are combined, a critical factor overlooked by previous methods (Georgiou et al., 2024).
+Therefore, we proposed the SASS strategy based on emotional semantic distance and screened the feature samples within each batch before mixing.
+Specifically, we use the cosine distance between each sample to represent semantic information similarity [38], and mix sample pairs with similarity greater than the threshold $\delta$ within each batch.
+This ensures that only samples with analogous emotions are combined, a critical factor that previous methods overlooked [17].
 We first performed $L_{2}$ normalization on the features of each modality $\textbf{Z}^{m}$ within the batch to ensure consistent feature scaling across modalities:
-(4) $\mathbf{Z}^{m}_{\text{norm}}=\frac{\mathbf{Z}^{m}}{\sqrt{\sum_{i=1}^{d^{m}}(Z_{i}^{m})^{2}}}.$
+$\mathbf{Z}^{m}_{\text{norm}}=\frac{\mathbf{Z}^{m}}{\sqrt{\sum_{i=1}^{d^{m}}(Z_{i}^{m})^{2}}}.$ (3)
 Then, we obtain the similarity matrix S as:
-(5) $\textbf{S}=\sum_{m\in t,v,a}{\mathbf{Z}^{m}_{\text{norm}}}\cdot(\mathbf{Z}^{m}_{\text{norm}})^{\top}/3.$
-Since the matrix S is symmetric ($s_{ij}=s_{ji},s\in\textbf{S}$) and the diagonal elements represent the self-similarity, we define the normalized upper triangular matrix $\textbf{S}^{\prime}$ as:
-(6) $\textbf{S}^{\prime}=\frac{1}{n(n-1)}\sum_{i\neq j}s_{ij}.$
-Then we randomly selected $B$ pairs to mix from the feature pairs $\{(\textbf{z}_{i},y_{i}),(\textbf{z}_{j},y_{j})\},\textbf{z}\in\textbf{Z}$, with a similarity greater than $\delta$ = 0.2.
+$\textbf{S}=\sum_{m\in t,v,a}{\mathbf{Z}^{m}_{\text{norm}}}\cdot(\mathbf{Z}^{m}_{\text{norm}})^{\top}/3.$ (4)
+Since the matrix S is symmetric ($s\in\textbf{S},s_{ij}=s_{ji}$) and the diagonal elements represent the self-similarity, we define the normalized upper triangular matrix $\textbf{S}^{\prime}$ as:
+$\textbf{S}^{\prime}=\frac{1}{n(n-1)}\sum_{i\neq j}s_{ij}.$ (5)
+We randomly selected $B$ pairs to mix from the feature pairs $\{(\textbf{z}_{i},y_{i}),(\textbf{z}_{j},y_{j})\},\textbf{z}\in\textbf{Z}$, with a similarity greater than $\delta$ = 0.2.
 Our SASS strategy employs the similarity threshold to exclude sample pairs with opposite emotions before mixing.
 A lower threshold effectively prevents semantic confusion in the mixed samples while maximizing data diversity.
-3.3. Sentiment Intensity Guided Mixing Module
-The mix ratio $\lambda$ is one of the most important hyperparameters in mixup-based methods, used to balance the degree of mixing between two or more samples.
+3.3 Sentiment Intensity Guided Mixing Module
+The mixing ratio $\lambda$ is one of the most important hyperparameters in mixup-based methods, controlling the degree of mixing between two or more samples.
 $\lambda$ is usually sampled from the $Beta(\alpha,\alpha)$ distribution.
-However, in recent years, some methods (Kim et al., 2020b; Liu et al., 2022; Qin et al., 2024a) have been able to automatically optimize the mixture proportions based on the sample states, thereby achieving better matching between the data and the label.
+However, in recent years, some methods [39, 30, 33] have automatically optimized mix ratios based on sample states, thereby achieving better alignment between the data and the labels.
 To ensure that samples with richer emotional semantics contribute more substantially to the mixing process, we propose the SIG mixing module.
-Specifically, for each modality, a Multi-Head self-Attention (MHA) (Vaswani et al., 2017) encoder is trained to predict emotional intensity values $\textbf{I}^{m}$.
+Specifically, for each modality, a Multi-Head self-Attention (MHA) [40] encoder is trained to predict emotional intensity values $\textbf{I}^{m}$.
 This encoder comprises an MHA layer, a residual connection, layer normalization, and a tanh activation function.
 The modality-specific emotional intensity predictor is denoted as $f_{\phi}^{m}$ (Eq.
-7-9).
-These predictions are then utilized as weighting factors to adjust the mixing ratios $\lambda^{m}$ during augmentation dynamically.
-(7) $head_{i}=\mathrm{Softmax}\left(\frac{\textbf{Z}^{m}\textbf{W}_{i}^{Q}\cdot\textbf{Z}^{m}({\textbf{W}_{i}^{K}})^{\top}}{\sqrt{d_{k}}}\right)\textbf{Z}^{m}\textbf{W}_{i}^{V},$
-(8) $\mathrm{MHA}(\textbf{Z}^{m})=\mathrm{LN}(\mathrm{Concat}(head_{1},head_{2},.,head_{h})\textbf{W}^{O}+\textbf{Z}^{m}),$
-(9) $\textbf{I}^{m}=f_{\phi}^{m}(\textbf{Z}^{m})=\mathrm{tanh}(\mathrm{GobalPool}(\mathrm{MHA}(\textbf{Z}^{m}))).$
-where $\textbf{W}^{\{Q,K,V\}}={\{\textbf{W}^{Q},\textbf{W}^{K},\textbf{W}^{V}\}}$ are learnable weight matrices that map the input features to queries $Q$, keys $K$, and values $V$, $d_{k}$ is the dimension of $K$, LNLN denotes Layer Normalization (Ba et al., 2016), the number of attention heads $h$ is set to 4 or 6, and $\textbf{W}^{O}$ is the output linear transformation matrix.
+6-8).
+These predictions are then used as weighting factors to adjust the mixing ratios $\lambda^{m}$ during augmentation dynamically.
+$head_{i}=\mathrm{Softmax}\left(\frac{\textbf{Z}^{m}\textbf{W}_{i}^{Q}\cdot\textbf{Z}^{m}({\textbf{W}_{i}^{K}})^{\top}}{\sqrt{d_{k}}}\right)\textbf{Z}^{m}\textbf{W}_{i}^{V},$ (6)
+$\mathrm{MHA}(\textbf{Z}^{m})$ $=\mathrm{LN}(\mathrm{Concat}(head_{1},head_{2},.,head_{h})\textbf{W}^{O}+\textbf{Z}^{m}),$ (7)
+$\textbf{I}^{m}=f_{\phi}^{m}(\textbf{Z}^{m})=\mathrm{tanh}(\mathrm{GobalPool}(\mathrm{MHA}(\textbf{Z}^{m}))).$ (8)
+where $\textbf{W}^{\{Q,K,V\}}={\{\textbf{W}^{Q},\textbf{W}^{K},\textbf{W}^{V}\}}$ are learnable weight matrices that map the input features to queries $Q$, keys $K$, and values $V$, $d_{k}$ is the dimension of $K$, LNLN denotes Layer Normalization [41], the number of attention heads $h$ is set to 4 or 6, and $\textbf{W}^{O}$ is the output linear transformation matrix.
 Based on the intensity of emotions $\textbf{I}^{m}$ in Eq.
-9, we perform Min-Max normalization on it, which can further calculate the intermediate mixing weights ${\bm{\omega}}^{m}_{i}$ for each feature pair $\{(\textbf{z}_{i},y_{i}),(\textbf{z}_{j},y_{j})\}$:
-(10) $\bm{\omega}_{i}^{m}=\frac{|\textbf{I}_{i}^{m}|-min(\textbf{I}^{m})}{max(\textbf{I}^{m})-min(\textbf{I}^{m})+\epsilon}.$
+8, we perform Min-Max normalization on it in modality, which can further calculate the intermediate mixing weights ${\bm{\omega}}^{m}_{i}$ for each feature pair $\{(\textbf{z}_{i},y_{i}),(\textbf{z}_{j},y_{j})\}$:
+$\bm{\omega}_{i}^{m}=\frac{|\textbf{I}_{i}^{m}|-min(\textbf{I}^{m})}{max(\textbf{I}^{m})-min(\textbf{I}^{m})+\epsilon},$ (9)
 where $\epsilon$ represents the minimum value that prevents division by zero and is set to $10^{-8}$.
 Due to the mixed strategy of convex combination, the adaptive mixing ratio $\lambda^{m}_{i,j}$ between the $\textbf{z}_{i}$ and $\textbf{z}_{j}$ is:
-(11) $\lambda^{m}_{i,j}=(\frac{\bm{\omega}_{i}^{m}}{\bm{\omega}_{i}^{m}+\bm{\omega}_{j}^{m}}+\lambda_{base})/2,$
+$\lambda^{m}_{i,j}=(\frac{\bm{\omega}_{i}^{m}}{\bm{\omega}_{i}^{m}+\bm{\omega}_{j}^{m}}+\lambda_{base})/2,$ (10)
 where the $\lambda_{base}$ sampled from the $\text{Beta}(\alpha,\alpha)$.
 Then, we calculate the average of the mixing ratios $\lambda^{m}$ for each modality to obtain the mixing ratio $\lambda^{L}_{i,j}$ of the labels:
-(12) $\lambda^{L}_{i,j}=\sum_{m\in t,v,a}\lambda_{i,j}^{m}/3.$
+$\lambda^{L}_{i,j}=\sum_{m\in t,v,a}\lambda_{i,j}^{m}/3.$ (11)
 Finally, we can get the mixed feature $\hat{\textbf{z}}^{m}$ and mixed label $\hat{y}$ using Eq.
-2-3.
-Unlike the previous Mixup-based studies (Zhang et al., 2018; Verma et al., 2019; Yun et al., 2019; Uddin et al., 2021), our proposed SIG module adaptively determines the mixing ratios between samples based on sentiment intensity, maintaining end-to-end trainability and allowing for continuous optimization throughout the training process.
-3.4. Sentiment Alignment Loss Function
-To enhance the accuracy of the emotion intensity predictor, we introduce the SAL as an additional regularization term to align the predicted sentiment distribution with the ground-truth labels.
-In our task, we seek the mapping from the data $x$ to its label $y$ modeled by network $f_{\theta}$, where the network parameters $\theta$ are optimized by minimizing a loss function $\mathcal{L}_{task}$ using Mean Squared Error (MSE):
-(13) $\min_{\theta}\mathcal{L}_{task}(f_{\theta}(\textbf{x}),y).$
-Since the emotional labels are continuous values (Zadeh et al., 2016, 2018b; Yu et al., 2020), we consider the mixup regression task.
-Therefore, we get the mapping between the mixed $\hat{\textbf{x}}$ and $\hat{y}$ by optimizing mixed MSE loss $\mathcal{L}_{mixMSE}$ (Eq.14).
+2.
+Unlike previous Mixup-based studies [13, 34, 15, 16], our proposed SIG module adaptively determines the mixing ratios between samples based on sentiment intensity, thereby maintaining end-to-end trainability and enabling continuous optimization throughout training.
+3.4 Sentiment Alignment Loss Function
+To enhance the accuracy of the emotion intensity predictor, we introduce the SAL as an additional regularization term to align the predicted distribution with the ground-truth labels.
+Formally, our goal is to learn a mapping from an input sample x to its corresponding label $y$ using a network $f_{\psi}\circ f_{\theta}$.
+The network parameters $\psi$ and $\theta$ are typically optimized by minimizing a task-specific loss $\mathcal{L}_{task}$:
+$\min_{\psi,\theta}\mathcal{L}_{task}(f_{\psi}(f_{\theta}(\textbf{x})),y).$ (12)
+Since the emotional labels are continuous values [42, 43, 44], we consider the mixup regression task.
+Therefore, we get the mapping between the mixed $\hat{\textbf{x}}$ and $\hat{y}$ by optimizing mixed Mean Squared Error (MSE) loss $\mathcal{L}_{mixMSE}$ (Eq.13).
 By minimizing $\mathcal{L}_{mixMSE}$ (Eq.14), the consistency between the mixed samples and labels is constrained.
-(14) $\mathcal{L}_{mixMSE}=\lambda^{L}_{i,j}\cdot MSE(f_{\theta}(\hat{\textbf{x}_{i}}),y_{i})+(1-\lambda^{L}_{i,j})\cdot MSE(f_{\theta}(\hat{\textbf{x}_{j}}),y_{j}).$
+$\mathcal{L}_{mixMSE}$ $=\lambda^{L}_{i,j}\cdot MSE(f_{\psi}(f_{\theta}(\hat{\textbf{x}_{i}})),y_{i})+(1-\lambda^{L}_{i,j})\cdot MSE(f_{\psi}(f_{\theta}(\hat{\textbf{x}_{j}})),y_{j}).$ (13)
 […]
-The KL divergence (Van Erven and Harremos, 2014) is commonly used to measure the amount of information lost when an approximate distribution is used to represent a true distribution.
+The KL divergence [45] is commonly used to measure the amount of information lost when an approximate distribution is used to represent a true distribution.
 A smaller KL divergence indicates that the two distributions are more similar.
@@
 Specifically, we convert $\textbf{I}^{m}$ and ground truth Y into probability distributions $\textbf{P}^{m}$ and $\textbf{P}^{L}$ (Eq.15), then calculate the KL divergence between them (Eq.16).
-Finally, we scale the KL divergence and incorporate it as an additional regularization term (Eq.18).
+Finally, we scale the KL divergence and incorporate it as an additional regularization term (Eq.
+17).
 The scale factor $\beta$ is $10^{3}$.
 This approach further promotes alignment across different modalities in emotion label prediction.
-(16) $\textbf{P}^{m}=\frac{exp(\textbf{I}^{m})}{\sum^{B}_{i=1}exp(\textbf{I}^{m}_{i})},\textbf{P}^{L}=\frac{exp(\textbf{Y})}{\sum^{B}_{i=1}exp(y_{i})},$
-(17) $\mathrm{KL}(\textbf{P}^{L}||\textbf{P}^{m})=\frac{1}{B}\sum^{B}_{i=1}\textbf{P}^{L}_{i}(log\textbf{P}^{L}_{i}-log\textbf{P}^{m}_{i}),$
-(18) $\mathcal{L}_{SAL}=\sum_{m\in{t,v,a}}\mathrm{KL}(\textbf{P}^{L}||\textbf{P}^{m})\cdot\beta.$
+$\textbf{P}^{m}_{i}=\frac{exp(\textbf{I}^{m})}{\sum^{B}_{i=1}exp(\textbf{I}^{m}_{i})},\textbf{P}^{L}_{i}=\frac{exp(\textbf{Y})}{\sum^{B}_{i=1}exp(y_{i})},$ (15)
+$\mathrm{KL}(\textbf{P}^{L}||\textbf{P}^{m})=\frac{1}{B}\sum^{B}_{i=1}\textbf{P}^{L}_{i}(log\textbf{P}^{L}_{i}-log\textbf{P}^{m}_{i}),$ (16)
+$\mathcal{L}_{SAL}=\sum_{m\in{t,v,a}}\mathrm{KL}(\textbf{P}^{L}||\textbf{P}^{m})\cdot\beta.$ (17)
 In summary, the total loss $\mathcal{L}_{total}$ for model optimization is expressed as a joint loss defined as follows:
-(19) $\mathcal{L}_{total}=\mathcal{L}_{task}+\xi_{1}\cdot\mathcal{L}_{mixMSE}+\xi_{2}\cdot\mathcal{L}_{SAL}.$
-where $\xi_{1}$ and $\xi_{2}$ are the weight hyperparameters used to regulate mixed MSE loss and SAL loss, respectively, with the effects of $\xi_{1}$ and $\xi_{2}$ to be evaluated in our ablation study.
-The SAL can serve as an auxiliary regularization term that not only facilitates the training of the modality-specific predictor $f_{\phi}^{m}$, but also promotes consistency across different modalities.
-4. Experiments
-To evaluate the proposed approach, we conducted comprehensive experiments on three public benchmark datasets for MSA: CMU-MOSI (MOSI) (Zadeh et al., 2016), CMU-MOSEI (MOSEI) (Zadeh et al., 2018b), and CH-SIMS (SIMS) (Yu et al., 2020).
-We compare our method against three representative Mixup-based methods (Verma et al., 2019; Venkataramanan et al., 2023; Georgiou et al., 2024) on six state-of-the-art backbones (Zadeh et al., 2017; Liu and Shen, 2018; Tsai et al., 2019; Hazarika et al., 2020; Zhang et al., 2023; Zhuang et al., 2024).
-4.1. Datasets
+$\mathcal{L}_{total}=\mathcal{L}_{task}+\xi_{1}\cdot\mathcal{L}_{mixMSE}+\xi_{2}\cdot\mathcal{L}_{SAL}.$ (18)
+where $\xi_{1}$ and $\xi_{2}$ are the weight hyperparameters used to regulate mixed MSE loss and SAL loss, respectively, with the effects to be evaluated in Section 4.6.
+The SAL can serve as an auxiliary regularization term that not only facilitates training of the modality-specific predictor $f_{\phi}^{m}$ but also promotes consistency across modalities.
+Figure 3: The t-SNE visualization of the original features and mixed features generated by MS-Mix (a) and $\mathcal{P}$owMix (b) on the MOSEI dataset using the MISA model.
+We employ a color scheme (blue/red) to differentiate the positive and negative categories, and use transparency to distinguish the original features from the mixed ones.
+4 Experiments
+To evaluate the proposed approach, we conducted comprehensive experiments on three public MSA datasets: CMU-MOSI (MOSI) [42], CMU-MOSEI (MOSEI) [43], and CH-SIMS (SIMS) [44].
+We compare our method against four representative Mixup-based methods [34, 18, 17, 33] on six state-of-the-art backbones [21, 22, 12, 46, 5, 6].
+The comparative results are summarized in Tables 1,2,3.
+In addition, we conducted extensive ablation experiments (Table 4) and parameter sensitivity analysis (Fig.
+5), as well as visualizations of the model performance (Fig.
+3 and 4).
+4.1 Datasets
 These datasets differ in scale, data collection methods, and linguistic characteristics, thereby offering diverse and challenging testbeds for evaluating model robustness and generalization.
@@
 MOSI
-The MOSI (Zadeh et al., 2016) dataset is an English benchmark for emotion recognition, consisting of 2,199 opinion segments from 93 YouTube movie reviews.
-It offers aligned text (transcripts), audio, and visual (facial) data, annotated for sentiment intensity on a continuous scale from -3 (strongly negative) to +3 (strongly positive) by five independent annotators, with final labels calculated from the average scores.
+The MOSI [42] dataset is an English benchmark for emotion recognition, consisting of 2,199 opinion segments from 93 YouTube movie reviews.
+It offers aligned text (transcripts), audio, and visual (facial) data, annotated with sentiment intensity on a continuous scale from -3 (strongly negative) to +3 (strongly positive) by five independent annotators, with final labels calculated as the average of the annotators’ scores.
 The dataset reflects real-world complexity, including modality incongruity (e.g., sarcasm), and contains about 2.2 hours of video from 89 speakers.
-Widely utilized in multimodal fusion research, this dataset encompasses pre-extracted features (e.g., acoustic data, facial action units) alongside speaker metadata.
+Widely used in multimodal fusion research, this dataset comprises pre-extracted features (e.g., acoustic data and facial action units) alongside speaker metadata.
 MOSEI
-The MOSEI (Zadeh et al., 2018b) dataset includes 23,453 video segments from more than 1,000 different speakers covering over 250 topics.
+The MOSEI [43] dataset includes 23,453 video segments from more than 1,000 different speakers covering over 250 topics.
 It is the largest multimodal benchmark for sentiment and emotion analysis.
-The dataset provides aligned text transcripts, audio signals, and visual frames, all annotated by trained annotators with continuous sentiment intensity scores from -3 to 3 like MOSI, and six emotion intensities (happiness, sadness, anger, fear, disgust, surprise).
+The dataset provides aligned text transcripts, audio signals, and visual frames, all annotated by trained annotators with continuous sentiment intensity scores ranging from -3 to 3 (e.g., MOSI) and six emotion intensities (happiness, sadness, anger, fear, disgust, surprise).
 It captures real-world communication nuances, including cross-modal conflicts, and supports research in multimodal fusion, emotion recognition, and sarcasm detection.
-Preprocessed features of MOSI and MOSEI are available through the CMU-Multimodal SDK (Zadeh et al., 2018c).
+Preprocessed features of MOSI and MOSEI are available through the CMU-Multimodal SDK [47].
 SIMS
-The SIMS (Yu et al., 2020) dataset is a Chinese multimodal benchmark featuring 2,281 video segments from 200 diverse online videos (vlogs, reviews, dialogues), designed to address real-world modality asynchrony.
-It provides aligned text transcripts, audio, and visual data with annotations of unified sentiment scores on a continuous scale from -1 (strongly negative) to +1 (strongly positive).
-As the first Chinese dataset emphasizing cross-modal dynamics, SIMS supports research in asynchronous multimodal fusion, cross-modal interaction modeling, and Mandarin sentiment analysis, with raw videos and preprocessed features publicly available.
-4.2. Backbones and Baseline Methods
-To ensure a comprehensive comparison, we evaluated the proposed method against three mixup-based methods: Manifold Mixup (Verma et al., 2019), MultiMix (Venkataramanan et al., 2023), and $\mathcal{P}ow$Mix (Georgiou et al., 2024) on six MSA architectures: TFN (Zadeh et al., 2017), LMF (Liu and Shen, 2018), MuIT (Tsai et al., 2019), MISA (Hazarika et al., 2020), ALMT (Zhang et al., 2023), and GLoMo (Zhuang et al., 2024).
-These architectures represent a diverse spectrum of feature extraction architectures and fusion strategies, covering both early and recent advances in the field, as summarized in Table 1.
-Table 1. Description of the six MSA backbones used in the paper
-Model TFN (Zadeh et al., 2017) LMF (Liu and Shen, 2018) MuIT (Tsai et al., 2019) MISA (Hazarika et al., 2020) ALMT (Zhang et al., 2023) GLoMo (Zhuang et al., 2024)
-Main encoder LSTM LSTM CNN LSTM Transformer Transformer
-Fusion method Tensor fusion Low-rank Fusion Cross-attention Self-attention Cross-attention Self-attention
-Published In, Year EMNLP, 2017 ACL, 2018 ACL, 2019 ACM MM, 2020 EMNLP, 2023 ACM MM, 2024
-4.3. Experimental Settings
+The SIMS [44] dataset is a Chinese multimodal benchmark featuring 2,281 video segments from 200 diverse online videos (vlogs, reviews, dialogues).
+It provides aligned text transcripts, audio, and visual data, with unified sentiment scores on a continuous scale from -1 (strongly negative) to +1 (strongly positive).
+As the first Chinese dataset to emphasize cross-modal dynamics, SIMS supports research on asynchronous multimodal fusion, cross-modal interaction modeling, and Mandarin sentiment analysis, with raw videos and preprocessed features publicly available.
+4.2 Backbones and Baseline Methods
+To enable a comprehensive and systematic evaluation, we compared the proposed method against four representative mixup-based augmentation approaches: Manifold Mixup [34] performs random mixing in the latent space to encourage smoother decision boundaries; MultiMix [18] is specifically designed for multimodal fusion through modality-specific mixing operations; $\mathcal{P}$owMix [17] introduces a weight-aware mixing strategy that dynamically adjusts the mixing ratios according to the importance of each modality.
+AdvMixUp [33] leverages adversarial training to generate sample-dependent hard mixed samples using feature-level masks.
+We extend this method to the MSA task by adapting the input from raw images to features and applying modality-specific mixing for each modality.
+These comparisons were conducted across six widely-used MSA architectures: TFN [21], LMF [22], MuIT [12], MISA [46], ALMT [5], and GLoMo [6].
+These backbones encompass a broad range of feature extraction paradigms and fusion strategies, spanning from early influential works to more recent state-of-the-art approaches in MSA.
+All models were implemented using available code and evaluated under identical settings using the M-SENA framework [4].
+4.3 Experimental Settings
 Evaluation Metrics
-We adopt the evaluation metrics established in the M-SENA framework (Mao et al., 2022) unified framework to ensure consistent and comparable results.
-These include binary accuracy ($ACC_{2}$) and F1-score, multiclass accuracy ($ACC_{5}$, $ACC_{7}$) and Mean Absolute Error (MAE).
+We use the evaluation metrics defined in the M-SENA unified framework [4] to ensure consistent, comparable results.
+These include binary accuracy ($ACC_{2}$) and F1-score (F1), multiclass accuracy ($ACC_{5}$, $ACC_{7}$) and Mean Absolute Error (MAE).
 Among these, MAE is a regression metric.
 Classification accuracy values are derived by mapping the regression scores to discrete sentiment categories.
-The prefix "w-" denotes without neutral-labeled samples.
-Furthermore, to facilitate intuitive and comprehensive comparison across metrics, we have separately presented the average values (Avg.) of all the classification metrics (ACC and F1-score).
+The prefix ”w-” denotes the absence of neutral-labeled samples.
+Furthermore, to facilitate intuitive and comprehensive comparisons across metrics, we have presented the average values (Avg.) for all ACC and F1 metrics separately.
 Experimental Details
-All experiments were conducted on a high-performance server equipped with an NVIDIA GeForce RTX 3090 Ti GPU.
+The experiments utilized high-performance servers equipped with an NVIDIA GeForce RTX 3090 Ti or an NVIDIA H20 GPU.
 The code was implemented in Python 3.10.18 using the PyTorch 2.7.1 framework with CUDA 12.6 support.
-For a fair comparison, all models were implemented using publicly available code, and their original experimental settings were retained.
-Most of the models were evaluated within the M-SENA framework (Mao et al., 2022), ensuring that all the comparison methods are conducted under consistent experimental conditions.
-In MS-Mix, the base mixing ratio $\lambda_{base}$ is sampled from a symmetric Beta distribution with $\alpha=2.0$ and the similarity threshold $\delta$ is fixed at 0.2.
+In MS-Mix, the base mixing ratio $\lambda_{base}$ is sampled from a symmetric Beta distribution with $\alpha=2.0$, and the similarity threshold $\delta$ is fixed at 0.2.
 The loss weights $\xi_{1}$ and $\xi_{2}$ are set to 0.7 and 0.5, respectively.
-Additionally, the number of attention heads $h$ = 4 for the MOSI and MOSEI datasets, and $h$ = 6 for the SIMS dataset.
-Table 2.
-Results of Various Approaches on The MOSI Dataset.
-Bold: Best performance.
-Underline: Second-best performance.
-$\dagger$}: result reported in (Mao et al., 2022).
-*: result reported in (Georgiou et al., 2024).
-"-": result was not reported in the original paper.
-MODEL w-ACC2 (%) ACC2(%) w-F1-score(%) F1-score(%) ACC5(%) ACC7(%) MAE $\downarrow$ Avg.
+Additionally, the number of attention heads $h$ is 4 for the MOSI and MOSEI and 6 for the SIMS dataset.
+Table 1: Results of Various Approaches on The MOSI Dataset (Mean ± std over 5 Independent Runs).
+Bold: Best performance. $\dagger$}: result reported in [4].
+[…]
 TFN 79.08 77.99 79.11 77.95 39.39 34.46 0.947 64.66
-+ Manifold Mix 79.27 78.13 79.25 78.06 40.96 35.71 0.927 65.23
-+ MultiMix 79.57 78.43 79.60 78.40 39.07 34.69 0.922 64.96
-+ $\mathcal{P}ow$Mix 79.27 78.13 77.96 79.15 41.11 36.73 0.919 65.39
-+ MS-Mix (ours) 80.18 78.72 80.05 78.52 41.25 34.84 0.904 65.59
++ Manifold Mix 79.26±1.07 78.13±1.12 79.07±1.04 78.15±1.11 39.02±1.15 35.23±0.74 0.921±0.012 64.81±0.95
++ MultiMix 79.43±0.63 78.22±1.51 79.31±1.05 78.03±1.12 39.48±0.71 34.95±0.26 0.925±0.013 64.90±0.57
++ AdvMixUp 78.48±2.56 78.13±0.38 79.50±1.17 77.74±1.41 39.58±0.70 34.98±0.63 0.923±0.006 64.73±0.51
++ $\mathcal{P}ow$Mix 79.64±1.42 78.18±0.83 79.35±0.96 77.95±0.97 40.11±1.08 36.16±0.44 0.919±0.006 65.23±0.31
++ MS-Mix (ours) 80.22±0.55 78.72±0.55 79.99±1.11 78.39±0.67 40.45±0.68 35.89±0.20 0.914±0.007 65.61±0.29
 LMF 79.18 77.90 79.15 77.80 38.13 33.82 0.950 64.33
-+ Manifold Mix 79.42 77.84 79.50 77.86 38.92 35.13 0.948 64.78
-+ MultiMix 79.88 78.72 79.96 78.74 38.80 36.32 0.907 65.40
-+ $\mathcal{P}ow$Mix 80.49 79.01 80.54 79.00 43.73 38.48 0.895 66.88
-+ MS-Mix (ours) 82.16 80.90 82.12 80.81 42.42 36.88 0.893 67.55
++ Manifold Mix 79.23±1.12 77.85±1.28 79.34±0.94 77.95±0.68 37.84±1.22 34.17±0.32 0.943±0.009 64.40±0.21
++ MultiMix 79.92±0.81 78.19±0.24 80.29±0.89 78.58±0.46 38.51±0.49 35.80±0.65 0.925±0.006 65.21±0.31
++ AdvMixUp 79.27±1.06 77.84±0.25 79.40±0.62 78.63±0.53 38.16±0.48 34.70±0.90 0.928±0.004 64.67±0.41
++ $\mathcal{P}ow$Mix 80.05±0.45 78.90±0.30 79.86±0.81 78.55±0.54 38.95±0.83 35.91±0.51 0.915±0.009 65.37±0.39
++ MS-Mix (ours) 82.02±0.26 79.16±0.32 82.01±0.16 79.27±0.68 41.90±0.42 36.52±0.50 0.879±0.003 66.81±0.13
 MuIT 80.98 79.71 80.95 79.63 42.68 36.91 0.878 66.81
-+ Manifold Mix 81.55 79.88 81.48 79.74 43.11 36.55 0.838 67.05
-+ MultiMix 81.64 79.01 81.66 79.96 43.38 36.57 0.859 67.04
-+ $\mathcal{P}ow$Mix* 81.01 - 80.99 - 40.65 35.00 0.904 -
-+ $\mathcal{P}ow$Mix 81.11 79.65 80.64 79.41 41.79 35.25 0.880 66.31
-+ MS-Mix (ours) 81.84 79.86 81.80 79.87 43.48 37.51 0.827 67.39
++ Manifold Mix 81.01±0.58 80.38±1.26 80.04±1.95 79.34±0.28 41.14±1.67 36.66±1.37 0.864±0.019 66.43±0.83
++ MultiMix 81.49±0.72 80.20±0.67 81.62±0.48 80.19±0.91 42.76±0.87 37.37±0.69 0.856±0.004 67.27±0.47
++ AdvMixUp 81.20±0.55 79.19±0.27 81.64±0.39 79.57±0.74 42.64±0.75 36.51±0.46 0.859±0.007 66.79±0.28
++ $\mathcal{P}ow$Mix 81.44±0.56 79.70±0.62 81.38±0.52 79.53±0.84 41.57±0.62 36.39±0.87 0.876±0.04 66.67±0.27
++ MS-Mix (ours) 82.02±0.28 79.69±0.66 81.97±0.40 79.89±0.84 43.33±0.67 37.48±0.66 0.830±0.006 67.40±0.11
 MISA 83.54 81.84 83.58 81.82 47.08 41.37 0.777 69.87
-+ Manifold Mix 83.69 82.22 83.69 82.17 46.50 41.11 0.765 69.90
-+ MultiMix 83.38 81.34 83.28 81.16 47.38 41.69 0.749 69.71
-+ $\mathcal{P}ow$Mix* 83.49 - 83.50 - 48.02 42.65 0.761 -
-+ $\mathcal{P}ow$Mix 83.27 81.33 83.41 81.42 47.91 41.87 0.763 69.87
-+ MS-Mix (ours) 83.99 82.36 83.96 82.28 48.25 41.98 0.746 70.47
-ALMT 84.60 82.65 84.58 82.57 51.75 46.21 0.723 72.06
-+ Manifold Mix 83.99 82.22 83.98 82.14 52.62 46.65 0.742 71.93
-+ MultiMix 83.69 82.36 83.68 82.30 49.27 44.02 0.751 70.08
-+ $\mathcal{P}ow$Mix* 84.30 - 84.30 - 49.24 44.25 0.741 -
-+ $\mathcal{P}ow$Mix 84.62 81.63 84.68 81.65 51.96 46.27 0.726 71.80
-+ MS-Mix (ours) 84.99 83.22 84.96 82.73 52.19 47.06 0.720 72.36
-GLoMo 84.89 83.50 84.92 83.94 53.87 46.13 0.749 72.88
-+ Manifold Mix 85.14 83.18 85.07 83.48 53.93 52.27 0.729 73.85
-+ MultiMix 84.70 83.42 84.60 83.64 53.82 52.19 0.743 73.73
-+ $\mathcal{P}ow$Mix 84.77 84.04 84.71 83.90 53.90 47.53 0.733 73.14
-+ MS-Mix (ours) 85.66 83.55 85.58 83.82 54.58 52.90 0.724 74.35
-Fig.
-3.
-The t-SNE visualization of the original features and mixed features generated by MS-Mix (a) and $\mathcal{P}$owMix (b) on the MOSEI dataset using the MISA model.
-We employ a color scheme (blue/red) to differentiate the positive and negative categories, and use transparency to distinguish the original features from the mixed ones.
-:
-4.4. Verification Performance of MS-Mix
-This chapter evaluates our proposed method, MS-Mix, on three publicly available MSA datasets: MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018b), and SIMS (Yu et al., 2020).
-Experiments were conducted using six backbone models, representing both classical (Zadeh et al., 2017; Liu and Shen, 2018; Tsai et al., 2019; Hazarika et al., 2020) and state-of-the-art approaches (Zhang et al., 2023; Zhuang et al., 2024) in terms of model architectures and fusion strategies, to demonstrate the general applicability of MS-Mix across different frameworks.
-For each model, we compared against classical Mixup variants: Manifold Mixup (Manifold Mix) (Verma et al., 2019) and MultiMix (Venkataramanan et al., 2023), as well as a recently proposed mixup-based method in MSA, $\mathcal{P}ow$Mix (Georgiou et al., 2024).
-The comparative results are summarized in Tables 2,3,4.
-Table 3.
-Results of Various Approaches on the MOSEI Dataset.
-Bold: Best performance.
-Underline: Second-best performance.
-$\dagger$}: result reported in (Mao et al., 2022).
-*: result reported in (Georgiou et al., 2024).
-"-": result was not reported in the original paper.
++ Manifold Mix 83.03±0.84 81.49±0.53 83.07±0.84 81.47±0.51 48.00±0.85 41.84±0.38 0.762±0.013 69.82±0.42
++ MultiMix 82.91±0.66 81.80±0.41 82.97±0.91 81.88±0.51 47.40±0.77 42.04±0.36 0.751±0.005 69.83±0.31
++ AdvMixUp 83.44±0.47 81.39±0.50 83.34±0.50 81.34±0.69 47.59±0.59 41.56±0.65 0.762±0.007 69.78±0.26
++ $\mathcal{P}ow$Mix 83.17±0.80 81.45±0.65 83.43±0.46 81.41±0.47 48.16±0.31 42.03±0.55 0.757±0.010 69.94±0.36
++ MS-Mix (ours) 83.94±0.52 82.21±0.63 83.81±0.32 82.14±0.37 48.59±0.56 42.36±0.31 0.746±0.005 70.51±0.19
+ALMT 83.48±0.80 82.06±0.61 83.39±0.78 81.94±0.59 51.32±0.75 45.46±0.86 0.733±0.003 71.28±0.23
++ Manifold Mix 84.16±0.76 82.58±0.82 83.58±0.50 82.60±0.77 51.51±0.93 46.04±1.39 0.734±0.008 71.74±0.15
++ MultiMix 83.53±0.71 82.05±0.38 83.31±1.03 81.89±0.61 48.92±0.74 44.21±0.65 0.747±0.005 70.65±0.38
++ AdvMixUp 83.95±0.23 81.94±0.35 83.61±0.67 81.95±0.68 51.44±1.20 46.89±0.59 0.726±0.004 71.63±0.21
++ $\mathcal{P}ow$Mix 83.38±0.50 82.34±0.96 83.26±0.84 82.33±0.93 51.67±0.53 45.72±0.64 0.730±0.007 71.45±0.25
++ MS-Mix (ours) 84.02±0.51 83.21±0.60 83.93±0.33 83.02±0.28 52.33±0.65 46.98±0.37 0.721±0.004 72.25±0.32
+GLoMo 83.81±0.27 82.19±0.32 83.73±0.28 82.13±0.33 52.59±0.71 47.10±0.94 0.743±0.010 71.92±0.22
++ Manifold Mix 84.78±0.43 82.95±0.62 84.80±0.23 83.00±0.70 52.89±0.96 49.45±0.99 0.731±0.006 72.98±0.18
++ MultiMix 84.94±0.56 83.23±0.58 85.02±0.49 83.32±0.58 53.01±0.47 50.46±0.93 0.728±0.004 73.33±0.22
++ AdvMixUp 84.27±0.73 83.24±0.65 84.39±0.59 83.25±0.65 53.27±0.49 51.09±0.68 0.730±0.004 73.25±0.36
++ $\mathcal{P}ow$Mix 84.81±0.29 83.36±0.19 84.84±0.36 83.48±0.35 53.14±0.75 49.56±2.06 0.730±0.007 73.20±0.31
++ MS-Mix (ours) 85.21±0.63 83.46±0.19 85.23±0.56 83.43±0.32 54.64±0.96 52.31±0.59 0.726±0.005 74.04±0.47
+4.4 Verification Performance of MS-Mix
+This chapter evaluates MS-Mix on three publicly available MSA datasets: MOSI [42], MOSEI [43], and SIMS [44].
+Experiments were conducted using six backbone models, representing both classical [21, 22, 12, 46] and state-of-the-art approaches [5, 6] in terms of model architectures and fusion strategies, to demonstrate the general applicability of MS-Mix across different frameworks.
+For each model, we compared against classical Mixup variants, including Manifold Mixup (Manifold Mix) [34] and Multimix [18], as well as a recently proposed mixup-based method in MSA, $\mathcal{P}ow$Mix [17], and the adversarial-based Mixup method, AdvMixUp [33].
+To enhance the reliability and robustness of our results, all experiments were conducted five times with different random seeds, and we report the mean and standard deviation (std) for each metric.
+As shown in Tables 1 to 3, MS-Mix achieves top performance, ranking first or second across all metrics compared with the baseline and other mixup-based methods.
+It is particularly noteworthy that MS-Mix consistently outperforms all competing approaches on the comprehensive Avg.
+metric across every backbone model.
+Table 2: Results of Various Approaches on the MOSEI Dataset (Mean ± std over 5 Independent Runs).
+Bold: Best performance. $\dagger$}: result reported in [4].
 […]
 TFN 81.89 78.50 81.74 78.96 53.10 51.60 0.573 70.97
-+ Manifold Mix 84.31 82.91 84.08 83.03 52.26 51.02 0.572 72.94
-+ MultiMix 82.94 80.47 82.80 80.82 53.51 52.39 0.568 72.16
-+ $\mathcal{P}ow$Mix 83.46 81.61 83.31 81.87 53.36 51.79 0.562 72.57
-+ MS-Mix (ours) 83.63 82.40 83.23 82.36 54.00 52.48 0.557 73.02
++ Manifold Mix 83.41±0.73 81.42±0.48 83.42±0.61 81.51±0.65 52.87±0.65 51.25±0.72 0.571±0.003 72.31±0.19
++ MultiMix 82.73±0.50 80.61±0.63 82.46±0.64 80.46±0.98 53.17±0.74 51.74±0.64 0.569±0.003 71.86±0.32
++ AdvMixUp 82.85±0.51 80.77±0.71 82.96±0.65 80.98±0.54 53.32±0.63 52.20±0.69 0.566±0.002 72.18±0.16
++ $\mathcal{P}ow$Mix 83.36±1.00 81.25±0.65 83.37±0.72 81.51±0.53 53.08±0.58 51.64±0.55 0.564±0.004 72.37±0.56
++ MS-Mix (ours) 83.41±0.39 82.16±0.26 83.30±0.49 82.07±0.31 53.87±0.44 52.20±0.29 0.559±0.002 72.84±0.17
 LMF 83.48 80.54 83.36 80.94 52.99 51.59 0.576 72.15
-+ Manifold Mix 83.74 78.71 83.85 78.98 53.53 52.33 0.563 71.86
-+ MultiMix 84.51 81.07 84.44 81.52 54.15 52.65 0.553 73.06
-+ $\mathcal{P}ow$Mix 83.38 80.30 83.33 80.79 54.30 53.10 0.559 72.53
-+ MS-Mix (ours) 84.65 83.30 84.28 83.26 54.13 52.74 0.554 73.73
++ Manifold Mix 83.21±0.64 78.91±1.49 83.39±0.68 79.01±1.46 53.40±0.83 52.04±0.76 0.572±0.008 71.66±0.44
++ MultiMix 84.17±0.71 80.73±0.80 84.29±0.58 80.80±0.83 53.72±0.88 52.72±0.96 0.564±0.002 72.74±0.36
++ AdvMixUp 84.12±0.55 81.72±0.81 84.12±0.63 81.79±0.75 53.47±1.04 52.14±0.77 0.565±0.008 72.90±0.38
++ $\mathcal{P}ow$Mix 83.95±0.31 80.97±0.59 84.02±0.30 81.27±0.53 54.37±0.59 52.73±0.74 0.560±0.006 72.88±0.30
++ MS-Mix (ours) 84.53±0.54 82.99±0.48 84.60±0.40 83.04±0.38 54.45±0.43 52.65±0.47 0.556±0.004 73.71±0.27
 MuIT 84.63 81.15 84.52 81.56 54.51 52.84 0.559 73.20
-+ Manifold Mix 84.46 81.64 84.54 81.40 54.02 52.51 0.561 73.10
-+ MultiMix 84.82 81.31 84.61 81.03 54.79 52.07 0.562 73.11
-+ $\mathcal{P}ow$Mix* 84.44 - 84.38 - 54.26 52.75 0.559 -
-+ $\mathcal{P}ow$Mix 84.37 81.45 84.56 81.34 54.57 52.45 0.560 73.12
-+ MS-Mix (ours) 84.68 81.69 84.64 82.07 54.98 53.09 0.556 73.53
++ Manifold Mix 84.39±0.27 81.31±1.19 84.32±0.25 81.55±1.01 53.94±0.48 52.39±0.48 0.562±0.002 72.98±0.58
++ MultiMix 83.99±0.36 81.45±0.68 84.03±0.53 81.75±0.61 54.22±0.52 52.46±0.49 0.565±0.006 72.98±0.19
++ AdvMixUp 84.31±0.50 81.17±0.17 84.29±0.71 80.94±0.74 54.30±0.44 51.98±0.17 0.570±0.003 72.83±0.33
++ $\mathcal{P}ow$Mix 84.71±0.62 81.74±0.75 84.65±0.90 81.85±0.35 54.37±0.51 52.68±0.47 0.559±0.003 73.33±0.23
++ MS-Mix (ours) 84.98±0.34 82.08±0.37 84.87±0.73 82.19±0.53 55.00±0.39 53.04±0.28 0.555±0.003 73.69±0.27
 MISA 84.67 80.67 84.66 81.12 53.63 52.05 0.558 72.80
-+ Manifold Mix 84.67 80.94 84.67 81.48 53.85 52.05 0.547 72.94
-+ MultiMix 84.78 80.49 84.79 81.08 53.68 52.29 0.544 72.85
-+ $\mathcal{P}ow$Mix* 84.97 - 84.86 - 54.52 53.00 0.543 -
-+ $\mathcal{P}ow$Mix 84.50 81.34 84.75 82.02 53.97 52.87 0.544 73.24
-+ MS-Mix (ours) 84.76 83.60 84.95 83.47 54.78 53.06 0.543 74.10
-ALMT 85.33 81.13 85.31 81.66 54.71 52.84 0.542 73.50
-+ Manifold Mix 86.54 83.69 86.37 83.90 53.38 51.51 0.542 74.23
-+ MultiMix 85.64 81.73 85.65 82.26 54.04 54.39 0.531 73.95
-+ $\mathcal{P}ow$Mix* 85.85 - 85.94 - 54.89 53.29 0.535 -
-+ $\mathcal{P}ow$Mix 85.77 81.25 85.42 81.92 54.17 52.89 0.541 73.57
-+ MS-Mix (ours) 85.91 83.22 85.78 83.48 55.18 53.60 0.540 74.53
-GLoMo 85.00 83.44 84.88 83.65 54.21 52.47 0.543 73.94
-+ Manifold Mix 85.94 83.85 85.80 84.05 55.16 53.09 0.546 74.65
-+ MultiMix 84.97 83.90 84.87 83.72 54.55 51.82 0.547 73.97
-+ $\mathcal{P}ow$Mix 85.41 83.22 85.29 83.47 55.12 53.18 0.541 74.28
-+ MS-Mix (ours) 85.84 84.08 85.72 84.13 55.41 53.27 0.533 74.74
-Table 4.
-Results of Various Approaches on The SIMS Dataset.
++ Manifold Mix 84.51±0.81 80.24±2.22 84.57±0.73 81.18±1.99 54.49±0.55 52.83±0.69 0.548±0.006 72.90±0.90
++ MultiMix 84.23±0.53 81.31±0.47 84.20±0.68 81.26±0.33 54.07±0.47 52.66±0.38 0.552±0.002 72.95±0.33
++ AdvMixUp 84.65±0.58 81.69±0.49 84.55±0.93 81.64±0.64 54.79±0.57 52.69±0.35 0.544±0.009 73.34±0.19
++ $\mathcal{P}ow$Mix 84.49±0.49 82.03±0.29 84.19±0.38 81.95±0.19 54.79±0.46 52.20±0.62 0.547±0.003 73.27±0.18
++ MS-Mix (ours) 84.84±0.31 83.15±0.36 84.72±0.33 83.08±0.51 55.04±0.34 52.95±0.43 0.542±0.004 73.96±0.28
+ALMT 85.15±0.80 81.35±0.56 85.04±0.48 81.53±0.50 54.11±0.29 52.73±0.32 0.550±0.003 73.32±0.14
++ Manifold Mix 85.13±1.46 82.16±1.06 85.08±1.33 81.65±2.73 54.33±0.83 52.65±1.00 0.543±0.005 73.33±1.19
++ MultiMix 85.10±1.02 82.00±0.73 85.22±0.60 81.96±0.78 54.70±0.74 52.70±0.70 0.538±0.007 73.61±0.50
++ AdvMixUp 85.22±0.56 81.99±0.32 85.05±0.64 82.03±0.42 54.46±0.54 52.87±0.30 0.544±0.008 73.60±0.32
++ $\mathcal{P}ow$Mix 85.55±0.46 82.37±0.45 85.48±0.51 82.35±0.26 54.58±0.69 53.13±0.45 0.539±0.004 73.91±0.23
++ MS-Mix (ours) 85.41±0.62 82.90±0.66 85.62±0.55 83.08±0.56 55.17±0.21 53.55±0.36 0.536±0.003 74.29±0.30
+GLoMo 84.93±0.80 82.71±0.56 85.28±0.50 82.95±0.77 53.80±0.48 52.67±0.60 0.542±0.004 73.72±0.52
++ Manifold Mix 84.88±0.90 83.65±0.57 85.13±0.72 83.28±0.50 56.08±0.84 53.37±1.53 0.535±0.011 74.40±0.48
++ MultiMix 85.13±0.39 83.29±0.67 85.25±0.39 82.92±0.81 55.05±0.28 52.93±0.33 0.535±0.005 74.09±0.27
++ AdvMixUp 85.24±0.40 83.20±0.39 84.92±0.42 82.98±0.35 55.00±0.39 53.04±0.36 0.537±0.004 74.06±0.22
++ $\mathcal{P}ow$Mix 85.08±0.20 83.19±0.23 85.12±0.26 83.38±0.48 55.08±0.29 53.07±0.27 0.533±0.003 74.15±0.07
++ MS-Mix (ours) 85.32±0.26 84.12±0.31 85.25±0.35 84.17±0.38 55.84±0.22 53.15±0.30 0.531±0.003 74.64±0.10
+Table 3: Results of Various Approaches on The SIMS Dataset (Mean ± std over 5 Independent Runs).
+Bold: Best performance. $\dagger$}: result reported in [4].
+MODEL ACC2(%) $\uparrow$ ACC3(%) $\uparrow$ ACC5(%) $\uparrow$ F1(%) $\uparrow$ MAE $\downarrow$ Avg(%) $\uparrow$.
+TFN 78.38 65.12 39.30 78.62 0.432 65.36
++ Manifold Mix 78.56±1.13 66.77±0.91 39.54±0.73 78.62±1.01 0.437±0.006 65.87±0.72
++ MultiMix 79.61±0.62 65.95±0.80 37.23±0.86 79.71±0.74 0.438±0.003 65.63±0.26
++ AdvMixUp 79.42±0.96 67.15±1.48 39.33±0.35 79.43±0.62 0.427±0.007 66.33±0.73
++ $\mathcal{P}ow$Mix 79.85±0.56 67.89±0.67 39.50±0.44 79.95±0.52 0.428±0.003 66.80±0.29
++ MS-Mix (ours) 80.66±0.47 68.39±0.72 40.74±0.52 80.71±0.73 0.423±0.005 67.62±0.39
+LMF 77.77 64.68 40.53 77.88 0.441 65.22
++ Manifold Mix 78.08±0.82 65.85±0.68 38.48±0.92 78.04±0.49 0.447±0.004 65.11±0.52
++ MultiMix 77.97±0.27 65.55±0.64 37.61±1.04 77.70±0.51 0.448±0.011 64.71±0.30
++ AdvMixUp 78.09±0.71 65.96±0.47 39.49±0.89 78.16±0.68 0.435±0.007 65.43±0.63
++ $\mathcal{P}ow$Mix 77.91±0.41 65.88±0.17 39.64±0.49 78.04±0.42 0.437±0.002 65.37±0.25
++ MS-Mix (ours) 78.12±0.24 65.80±0.30 41.02±0.82 78.16±0.37 0.433±0.003 65.77±0.20
+MuIT 77.46 65.43 34.79 77.57 0.446 63.81
++ Manifold Mix 78.34±0.44 64.33±0.95 34.94±2.44 78.00±0.78 0.448±0.001 63.90±0.58
++ MultiMix 77.70±1.05 64.80±0.63 36.55±0.99 77.70±0.73 0.447±0.005 64.19±0.27
++ AdvMixUp 78.10±0.54 64.27±0.71 37.61±0.51 78.20±0.47 0.445±0.006 64.55±0.50
++ $\mathcal{P}ow$Mix 78.29±1.03 65.07±0.40 37.74±0.32 78.11±1.06 0.443±0.007 64.80±0.42
++ MS-Mix (ours) 78.51±0.71 65.53±0.43 38.16±0.18 78.70±1.13 0.443±0.003 65.22±0.48
+MISA 76.53±0.82 62.96±0.29 36.01±0.64 76.71±0.90 0.452±0.004 63.05±0.44
++ Manifold Mix 76.87±0.69 63.96±0.35 37.67±0.81 77.08±0.69 0.458±0.005 63.89±0.28
++ MultiMix 77.34±0.36 64.26±0.05 37.30±1.18 77.07±1.32 0.439±0.009 63.99±0.21
++ AdvMixUp 77.01±0.80 63.82±0.48 36.47±1.02 77.22±0.52 0.450±0.010 63.63±0.40
++ $\mathcal{P}ow$Mix 77.56±0.39 64.32±0.39 37.21±0.30 77.48±0.33 0.441±0.003 64.14±0.27
++ MS-Mix (ours) 78.19±0.28 64.70±0.59 37.75±0.21 78.14±0.44 0.434±0.004 64.69±0.23
+ALMT 77.19±0.88 64.25±0.80 38.68±0.50 76.81±0.86 0.442±0.009 64.23±0.48
++ Manifold Mix 76.88±1.49 63.67±1.00 38.49±1.94 77.08±1.22 0.429±0.004 64.03±0.49
++ MultiMix 77.52±1.30 63.04±0.42 38.62±0.52 78.11±0.09 0.436±0.007 64.32±0.17
++ AdvMixUp 77.84±1.14 64.50±1.30 37.94±0.30 77.88±0.64 0.427±0.003 64.54±0.08
++ $\mathcal{P}ow$Mix 77.61±0.25 64.84±0.77 38.44±0.61 77.87±0.34 0.424±0.004 64.69±0.36
++ MS-Mix (ours) 78.21±0.14 64.52±0.53 38.63±0.41 78.30±0.40 0.426±0.003 64.92±0.36
+GLoMo 78.80±0.66 65.66±0.76 37.11±0.86 78.81±0.51 0.427±0.005 65.09±0.51
++ Manifold Mix 78.98±1.16 65.93±0.62 37.78±0.52 78.94±0.67 0.426±0.005 65.41±0.47
++ MultiMix 78.64±0.66 66.42±0.42 38.28±0.41 79.05±0.88 0.428±0.010 65.60±0.28
++ AdvMixUp 78.31±0.88 66.98±0.83 38.30±0.20 77.23±0.36 0.432±0.002 65.20±0.44
++ $\mathcal{P}ow$Mix 79.34±0.65 65.11±1.94 38.33±0.42 79.12±0.18 0.428±0.007 65.47±0.45
++ MS-Mix (ours) 79.33±0.45 67.38±0.98 38.36±0.31 79.23±0.30 0.423±0.001 66.07±0.39
+Figure 4: Visualization of loss landscapes comparing Mixup variants and MS-Mix on MOSEI using ALMT.
+The consistent outperformance of MS-Mix across diverse backbones and datasets stems from its core design principles, which directly address key limitations in existing mixup-based augmentation methods.
+Unlike these methods, which often rely on random or offline mixing strategies, MS-Mix incorporates three interconnected mechanisms tailored for MSA: First, the SASS strategy prevents the blending of semantically contradictory samples (e.g., happy and angry) by filtering pairs based on emotional similarity in the latent space.
+In contrast, the comparative methods [34, 18, 17, 33] mix random sample pairs regardless of emotional coherence, often yielding ambiguous or noisy samples that mislead the training process.
+By ensuring semantic and label consistency, SASS significantly enhances the quality of augmented data.
+Second, the SIG mixing module adaptively determines modality-specific mixing ratios based on emotional salience, rather than using random or fixed values as in Manifold Mix [34], MultiMix [18], or leveraging modality importance without emotional context like $\mathcal{P}ow$Mix [17] and AdvMixUp [33].
+This allows MS-Mix to dynamically weight each modality’s contribution, leading to more discriminative mixed features and improved robustness of learned cross-modal representations.
+Third, the SAL serves as a regularization term that aligns predicted sentiment distributions with ground-truth labels, enhancing prediction consistency across modalities and providing additional supervisory signal.
+This component is unique to MS-Mix and offers a critical advantage over comparative methods, which lack explicit constraints on the alignment of emotion-level distributions.
+As a result, SAL not only stabilizes training but also strengthens the model’s resistance to overfitting.
+Table 4: Results of the ablation experiments using the LMF method on the MOSI database.
 Bold: Best performance.
-Underline: Second-best performance.
-$\dagger$}: result reported in (Mao et al., 2022).
-*: result reported in (Georgiou et al., 2024).
-"-": result was not reported in the original paper.
-MODEL ACC2 ACC3 ACC5 F1-score MAE $\downarrow$ Avg.
-TFN 78.38 65.12 39.30 78.62 0.432 65.36
-+ Manifold Mix 79.21 67.83 40.48 78.92 0.443 66.61
-+ MultiMix 78.99 66.08 35.45 78.01 0.443 64.63
-+ $\mathcal{P}ow$Mix 79.21 67.40 37.42 79.06 0.423 65.77
-+ MS-Mix (ours) 81.18 68.71 40.04 81.14 0.421 67.77
-LMF 77.77 64.68 40.53 77.88 0.441 65.22
-+ Manifold Mix 77.90 65.86 36.98 76.90 0.453 64.41
-+ MultiMix 76.59 64.55 36.32 76.70 0.444 63.54
-+ $\mathcal{P}ow$Mix 77.86 66.96 39.54 77.33 0.441 65.42
-+ MS-Mix (ours) 78.34 65.65 40.61 77.92 0.433 65.63
-MuIT 77.46 65.43 34.79 77.57 0.446 63.81
-+ Manifold Mix 78.56 65.21 37.86 78.72 0.442 65.09
-+ MultiMix 78.24 65.21 38.11 78.41 0.447 64.99
-+ $\mathcal{P}ow$Mix* 79.04 - - 78.51 0.437 -
-+ $\mathcal{P}ow$Mix 78.64 65.49 37.85 78.90 0.444 65.22
-+ MS-Mix (ours) 78.77 65.79 38.73 78.62 0.441 65.48
-MISA 76.81 63.24 37.42 76.39 0.467 63.47
-+ Manifold Mix 77.19 64.01 36.29 77.05 0.456 63.64
-+ MultiMix 77.50 64.18 37.06 77.18 0.437 63.98
-+ $\mathcal{P}ow$Mix* 77.35 - - 76.97 0.441 -
-+ $\mathcal{P}ow$Mix 77.53 64.03 37.51 77.16 0.439 64.06
-+ MS-Mix (ours) 78.54 64.16 37.83 78.18 0.431 64.68
-ALMT 76.81 63.02 34.35 76.81 0.449 62.75
-+ Manifold Mix 76.37 64.99 36.32 76.66 0.460 63.59
-+ MultiMix 77.68 62.63 38.95 78.03 0.433 64.32
-+ $\mathcal{P}ow$Mix* 78.91 - - 79.13 0.429 -
-+ $\mathcal{P}ow$Mix 78.17 63.06 37.53 78.24 0.431 64.28
-+ MS-Mix (ours) 78.34 64.46 37.68 78.21 0.426 64.67
-GLoMo 79.09 66.31 37.75 79.20 0.427 65.59
-+ Manifold Mix 78.82 66.42 38.02 79.06 0.426 65.58
-+ MultiMix 78.89 66.38 38.40 79.27 0.424 65.74
-+ $\mathcal{P}ow$Mix 79.03 65.43 38.53 79.16 0.434 65.53
-+ MS-Mix (ours) 79.28 67.69 38.06 79.52 0.422 66.14
-As shown in Tables 2 to 4, MS-Mix achieves top performance, ranking first or second in all metrics compared to baseline and other mixup-based methods.
-It is particularly noteworthy that MS-Mix consistently outperforms all competing approaches on the comprehensive average (Avg.) metric across every backbone model.
-The consistent outperformance of MS-Mix across diverse backbones and datasets stems from its core design principles, which directly address key limitations in existing mixup-based augmentation methods such as Manifold Mixup (Verma et al., 2019), MultiMix (Venkataramanan et al., 2023), and $\mathcal{P}ow$Mix (Georgiou et al., 2024).
-Unlike these methods, which often rely on random or offline mixing strategies, MS-Mix incorporates three interconnected mechanisms tailored for MSA: First, the SASS strategy prevents the blending of semantically contradictory samples (e.g., happy and angry) by filtering pairs based on emotional similarity in the latent space.
-In contrast, the comparative methods perform mixing between random sample pairs regardless of emotional coherence, often resulting in ambiguous or noisy samples that mislead training.
-By ensuring semantic and label consistency, SASS significantly enhances the quality of augmented data.
-Second, the SIG mixing module adaptively determines modality-specific mixing ratios based on emotional salience, rather than using random or fixed values as in Manifold Mixup (Verma et al., 2019), MultiMix (Venkataramanan et al., 2023), or leveraging modality importance without emotional context like $\mathcal{P}ow$Mix.
-This allows MS-Mix to dynamically weight the contribution of each modality, leading to more discriminative mixed features and improved robustness of learned cross-modal representations.
-Third, the SAL serves as a regularization term that aligns predicted sentiment distributions with ground-truth labels, enhancing prediction consistency across modalities and providing additional supervisory signal.
-This component is unique to MS-Mix and offers a critical advantage over comparative methods, which lack explicit constraints on emotion-level distribution alignment.
-As a result, SAL not only stabilizes training but also strengthens the model’s resistance to overfitting.
-To further demonstrate the effectiveness of MS-Mix, as shown in Fig.
-3, we employed t-SNE (Maaten and Hinton, 2008) to visualize the textual modal feature distributions of both original and mixed features on the MOSEI dataset, using the MISA model as the backbone.
-The t-SNE visualization demonstrates that features generated by MS-Mix exhibit significantly clearer decision boundaries and more distinct cluster separation compared to those produced by $\mathcal{P}$owMix.
-This improvement indicates that MS-Mix not only preserves the semantic integrity of the original feature space but also generates more discriminative intermediate representations.
-Overall, MS-Mix effectively addresses the semantic confusion and label mismatch issues that hindered previous methods.
-By integrating three core strategies, our method generates discriminative features with corresponding labels, significantly enhancing the overall quality of augmented data.
-4.5. Ablation and Hyperparameter Sensitivity Analysis
-To evaluate the impact of key hyperparameters and individual algorithmic components of MS-Mix, we performed ablation studies on three core components using the LMF (Liu and Shen, 2018) model on the MOSI (Zadeh et al., 2016) and SIMS (Yu et al., 2020) datasets.
-Additionally, hyperparameter sensitivity analyses on the MOSEI dataset (Zadeh et al., 2018b), using the MISA (Hazarika et al., 2020) and GLoMo (Zhuang et al., 2024) models.
-In our ablation study, we integrated Manifold Mixup (Verma et al., 2019) into the classic LMF (Liu and Shen, 2018) backbone as the baseline to evaluate the contribution of each proposed component: SASS, SIG, and SAL.
-Each module was selectively enabled/disabled during training to assess its individual impact on the performance of MS-Mix.
-Note that since the computation of SAL relies on the output of the SIG module, it cannot be used independently.
-Results presented in Table 5 show that each of the three components contributes noticeably to performance gains over the baseline, underscoring the importance of every module.
-Furthermore, by incrementally incorporating these components, we observed progressive improvements in performance, demonstrating not only the individual efficacy of each module but also the significance of synergistic work.
-Table 5. Results of the ablation experiments using the LMF method on the MOSI database. Bold: Best performance.
 […]
-79.42 77.84 79.50 77.86 38.92 35.13 0.948 64.78
+79.23 77.85 79.34 77.95 37.84 34.17 0.943 64.40
 79.98 78.41 80.08 78.60 38.86 35.26 0.921 65.20
@@
 81.41 78.03 81.41 79.61 40.48 36.00 0.919 66.16
-82.16 80.90 82.12 80.81 42.42 36.88 0.893 67.55
-Fig.
-4.
-The impact of different parameter values.
+82.02 79.16 82.01 79.27 41.90 36.52 0.879 66.81
+SIG → Ground-Truth 79.76 78.29 79.82 78.31 38.27 34.87 0.936 64.89
+SIG → Linear Mapping 80.50 78.23 80.79 78.37 39.04 34.70 0.927 65.27
+4.5 Performance Visualization
+To further demonstrate the effectiveness of MS-Mix, we conduct both t-SNE [48] visualizations (Fig.
+3) and loss landscape visualizations (Fig.
+4).
+Specifically, we employ t-SNE to project the distributions of textual features for both the original and mixed features extracted from the MOSEI dataset, using the MISA model.
+The results reveal that features generated by MS-Mix exhibit significantly clearer decision boundaries and more distinct cluster separation than those produced by $\mathcal{P}$owMix.
+This enhanced separation indicates that MS-Mix not only preserves the semantic integrity of the original feature space but also yields more discriminative intermediate representations.
+Furthermore, we analyze the loss landscape to examine the generalization effect of MS-Mix.
+The generalization ability of neural networks is closely related to the geometry of their loss landscapes, where convergence to flat minima is generally associated with superior robustness and generalization performance [49].
+To visualize this, we plot the 2D loss surface of the ALMT model on MOSEI by perturbing the parameters along two random, mutually orthogonal directions, generated via Gaussian sampling and Gram-Schmidt orthogonalization.
+As shown in Fig.
+4, introducing Manifold Mix and other variants resulted in some landscape smoothing.
+However, pronounced peaks and high local curvature remained, indicating persistent sensitivity to parameter perturbations.
+In contrast, MS-Mix yields a notably smoother, flatter landscape, suggesting reduced gradient conflict during training and convergence toward more robust minima.
+In summary, by integrating the evidence from label distribution on MOSEI (Fig.
+1), feature visualization (Fig.
+3), and loss landscapes (Fig.
+4), MS-Mix effectively mitigates the semantic confusion and label mismatch issues inherent in other mixup variants.
+Our approach generates discriminative features with well-matched labels, thereby significantly enhancing the quality of mixed samples and improving overall performance in MSA.
+4.6 Ablation and Hyperparameter Analysis
+To evaluate the impact of key hyperparameters and individual algorithmic components of MS-Mix, we performed ablation studies on three core components using the LMF [22] model on the MOSI [42] and SIMS [44] datasets.
+Additionally, hyperparameter sensitivity analyses on the MOSEI dataset [43] were conducted using the MISA [46] and GLoMo [6] models.
+All experiments were run five times using different random seeds, and the average values were reported as the results.
+In our ablation study, we integrated Manifold Mix [34] into the classic LMF [22] backbone as the baseline to evaluate the contribution of each proposed component: SASS, SIG, and SAL.
+To isolate the effect of each module, we conducted experiments by selectively enabling or disabling individual components during training.
+Note that SAL relies on SIG’s output and thus cannot be evaluated independently.
+Furthermore, to justify the rationality of SIG’s complex design, we additionally investigated two simplified variants: degenerating SIG to a linear layer (SIG → Linear Mapping) and using the absolute values of the ground-truth labels as the intensity scores directly (SIG → Ground-Truth).
+All results were obtained by averaging over five independent runs with different random seeds.
+As shown in the table 4, each of the three core components contributes substantially to the overall performance gains over the baseline.
+Removing any single module results in noticeable degradation, underscoring its individual importance.
+Furthermore, simplified variants of SIG also lead to clear performance drops.
+This is because a simple linear mapping is insufficient for extracting the complex, non-linear nature of emotional intensity, and the direct use of a global label overlooks the differences in emotional expression across modalities.
+In contrast, the attention-based prediction mechanism in our full SIG module effectively captures these latent, modality-specific intensity signals.
+Figure 5: The impact of different parameter values.
 (a) Accuracy under different similarity thresholds $\delta$.
@@
 (d) Accuracy under different attention heads $h$.
-:
-To comprehensively analyze the sensitivity of key hyperparameters, we conducted extensive experiments using both the top-performing GLoMo model (Zhuang et al., 2024) and the classical MISA architecture (Hazarika et al., 2020) on the largest English dataset MOSEI (Zadeh et al., 2018b) and the Chinese benchmark SIMS (Yu et al., 2020).
+To comprehensively analyze the sensitivity of key hyperparameters, we conducted extensive experiments using both the top-performing GLoMo model [6] and the classical MISA architecture [46] on the largest English dataset MOSEI [43] and the Chinese benchmark SIMS [44].
 As illustrated in Fig.
-4, our evaluation reveals how critical hyperparameters affect model performance.
+5, our evaluation shows that hyperparameters are critical to model performance.
 Using the MISA model on MOSEI, we first examined the similarity selection threshold $\delta$ (Fig.
-4(a)).
+5(a)).
 The results demonstrate that at $\delta=0.2$, the method maintains an optimal balance, incorporating sufficient samples while effectively filtering out emotionally contradictory ones, thus achieving peak performance.
-Notably, even under suboptimal $\delta$ configurations, MS-Mix consistently surpasses Manifold Mixup (Verma et al., 2019).
+Notably, even under suboptimal $\delta$ configurations, MS-Mix consistently surpasses Manifold Mix [34].
 We further evaluated the loss weights $\xi_{1}$ and $\xi_{2}$ with the GLoMo model on MOSEI (Fig.
-4(b)).
+5(b)).
 The experimental outcomes confirm the rationality of our weight selection, showing that $\xi_{1}=0.7$ and $\xi_{2}=0.5$ yield the best performance.
 Additionally, we assessed the shape parameter $\alpha$ of the Beta distribution using MISA on MOSEI (Fig.
-4(c)), finding $\alpha=2.0$ to be optimal.
+5(c)), finding $\alpha=2.0$ to be optimal.
 The influence of the number of attention heads $h$ was also investigated across datasets (Fig.
-4(d)): $h=4$ delivers the best results on MOSEI, while $h=6$ is most effective on SIMS.
+5(d)): $h=4$ delivers the best results on MOSEI, while $h=6$ is most effective on SIMS.
 These findings collectively validate the rationale behind our parameter configurations.
-4.6. Occlusion Experiment
-In the random occlusion experiment, we compared the performance of our method against baseline approaches (Verma et al., 2019; Venkataramanan et al., 2023; Georgiou et al., 2024) under both clean and noisy data conditions across MOSI (Zadeh et al., 2016), MOSEI (Zadeh et al., 2018b), and SIMS (Yu et al., 2020).
-Specifically, we randomly masked out 0% to 40% of the input data during training of the LMF (Liu and Shen, 2018) model to investigate the effect of mixup-based augmentation on model robustness.
-Fig. 5. The performance of different mixup-based methods at different occlusion ratios on three datasets. :
-As illustrated in Fig.
-5, the experimental results demonstrate that the introduction of noise influences model performance across all datasets.
-A noteworthy phenomenon emerges in which moderate levels of random masking (e.g., 10% or 20%) unexpectedly improve performance, rather than degrade it.
+4.7 Occlusion Experiment
+In the random occlusion experiment, we compared the performance of MS-Mix against baseline methods [34, 18, 17] under both clean and noisy data conditions across MOSI [42], MOSEI [43], and SIMS [44].
+Specifically, we randomly masked out 0% to 40% of the input data during training of the LMF [22] model to investigate the effect of mixup-based augmentation on model robustness.
+Figure 6: The performance of different mixup-based methods at different occlusion ratios on three datasets.
+As shown in Fig.
+6, the experimental results indicate that introducing noise affects performance across all datasets.
+A noteworthy phenomenon emerges: moderate levels of random masking (e.g., 10% or 20%) unexpectedly improve performance rather than degrade it.
 Beyond these optimal masking ratios, however, model performance deteriorates.
-Despite these variations, MS-Mix consistently achieves the highest average accuracy across all evaluated conditions, demonstrating its advantages in optimizing the model for handling noisy data.
-5. CONCLUSION
-In this paper, we propose MS-Mix, a novel and adaptive data augmentation framework specifically designed for MSA.
+Despite these variations, MS-Mix consistently achieves the highest average accuracy across all evaluated conditions, demonstrating its advantage in optimizing the model to handle noisy data.
+5 CONCLUSION
+In this paper, we propose MS-Mix, a novel, adaptive data augmentation framework designed specifically for MSA.
 Unlike traditional mixup-based methods, which often rely on random and offline mixing strategies, MS-Mix introduces three key emotion-aware mechanisms to enhance the semantic consistency of augmented samples.
-First, the SASS module ensures that only emotionally consistent samples are mixed, effectively avoiding mixing feature pairs with contradictory emotions.
-Second, the SIG mixing module adaptively determines mixing ratios of each modality based on emotional salience, promoting the generation of samples with clearer classification boundaries.
-Third, the SAL aligns the predicted emotion distributions with ground-truth labels through the KL-divergence minimization, serving as an effective regularizer to improve generalization.
-Empirical evaluation of three benchmark datasets and six diverse backbone architectures has shown that MS-Mix achieves state-of-the-art performance.
-In the future, we will consider extending MS-Mix to other multimodal tasks, such as mental health monitoring and human-computer interaction, which could leverage its ability to maintain semantic consistency across a wider range of applications.
-We also intend to integrate self-supervised or semi-supervised learning strategies into MS-Mix to alleviate the reliance on large annotated datasets and enhance its applicability in low-resource settings.
-Finally, from an efficiency perspective, we will focus on developing lightweight variants of the sample selection and fusion mechanisms to reduce computational costs and support deployment on resource-constrained devices.
-6. Acknowledgments
-This work was supported by the Science and Technology Innovation Key R&D Program of Chongqing (No.
-CSTB2024TIAD-STX0023, CSTB2023TIAD-STX0015, and CSTB2023TIAD-STX0031), and in part by the Natural Science Foundation of China (No.
-62372427).
+First, the SASS module ensures that only emotionally consistent samples are combined, effectively avoiding the mixing of feature pairs with contradictory emotions.
+Second, the SIG mixing module adaptively determines mixing ratios for each modality based on emotional salience, thereby promoting samples with clearer decision boundaries.
+Third, the SAL aligns the predicted emotion distributions with ground-truth labels via KL-divergence minimization, serving as an effective regularizer that improves generalization.
+An empirical evaluation across three benchmark datasets and six diverse backbone architectures shows that MS-Mix achieves state-of-the-art performance.
+From an engineering perspective, MS-Mix enables more reliable and data-efficient multimodal emotion understanding, facilitating practical deployment in human-computer interaction, intelligent human-robot systems, and emotion-aware assistive technologies.
+In future work, we will consider extending MS-Mix to other multimodal tasks, such as mental health monitoring and human-computer interaction, leveraging its ability to maintain semantic consistency in a wider range of applications.
+Furthermore, we also intend to integrate self-supervised or semi-supervised learning strategies into MS-Mix to reduce reliance on large annotated datasets.
+6 Conflicts of Interest Statement
+The authors declare no conflicts of interest.